### v1

# Gradient Existence and Energy Finiteness of Local Minimizers in the Wasserstein $L^{\infty}$ Topology for Binary-Star Systems

Abstract
In this paper, we refine and complement McCann’s results on binary-star systems [31], a compressible fluid model governed by the Euler-Poisson equations. We consider a general form of the equation of state that includes polytropic gaseous stars indexed by $\gamma$ as a special case. Beyond revisiting McCann’s framework and conclusions—where solutions to the Euler-Poisson equations are obtained as local energy minimizers via variational methods under the topology induced by the Wasserstein $L^{\infty}$ distance—we focus on a detailed exploration of the properties of local energy minimizers in this topology, addressing three key aspects: (1) the feasibility of transitioning from the Euler-Lagrange equation to the Euler-Poisson equation by demonstrating gradient existence; (2) the existence of $L^{\infty}$ functions within neighborhoods in this topology; and (3) the finiteness of the energy of local minimizers in this topology, contrasted with the non-existence of finite-energy local minimizers and the existence of infinite-energy weak local minimizers in the topology inherited from topological vector spaces.
Key words: Gaseous Binary-Star systems, Euler-Poisson equations, Calculus of variations, Wasserstein-$L^{\infty}$ topology
1 Introduction
A fundamental hydrodynamic model for Newtonian stellar dynamics is given by the Euler–Poisson system, describing the evolution of an isolated self‑gravitating fluid.
$\begin{split}\partial_{t}\rho+\nabla\cdot\left({\rho v}\right)&=0,\\ \rho\partial_{t}v+\rho\left({v\cdot\nabla}\right)v+\nabla P(\rho)&=\rho\nabla V,\\ \Delta V&=-4\pi\rho.\end{split}$ (EP)
where $\rho(x,t)\geq 0$ is the density of fluids at position $x\in\mathbb{R}^{3}$ and time $t\geq 0$, $v(x,t)\in\mathbb{R}^{3}$ is the velocity, and $V(x,t)\in\mathbb{R}$ is the gravitational potential. Pressure $P(\rho)$ depends on the density only, with some appropriate assumptions similar to McCann’s [31].
The existence of single-star and multi-star systems has been established by variational methods (see [4, 5, 3, 10, 11, 26, 31]) and perturbative approaches (see [21, 22, 19, 36, 37, 1, 27]).
As Jang and Seok discussed in [23], in the case of a uniformly rotating system, where we simply consider only the orbital revolution around each other and ignore the rotations of objects around their own axes, the Euler–Poisson equations reduce to a single equation:
$\omega^{2}\tilde{\rho}\left(x\right)P_{12}\left(x\right)+\nabla P\left(\tilde{\rho}\left(x\right)\right)-\tilde{\rho}\left(x\right)\left(\nabla V_{\tilde{\rho}}\left(x\right)\right)=0$ (EP’)
where $\omega>0$ is angular velocity, $P_{12}\left(x_{1},x_{2},x_{3}\right)=(x_{1},x_{2},0)$ , $\tilde{\rho}$ is a density function (in general with compact support), and $V_{\tilde{\rho}}(x)={\int{\frac{\tilde{\rho}(y)}{\left|{y-x}\right|}d^{3}y}}$. And the discussions in subsection 2.2 below shows that once $\tilde{\rho}$ satisfies the reduced Euler-Poisson equation (EP’), then $\rho\left({t,y}\right)=\tilde{\rho}\left(R_{-\omega t}y\right)$, $v\left({t,y}\right)=\omega\left({-y_{2},y_{1},0}\right)^{T}$ satisfies the original system (EP), where $R_{-\omega t}y$ is the rotation map given in Definition 2.1.
In [31], McCann used a variational method to construct uniformly rotating binary‑star solutions. His binary stars exhibit a dual characterization: as Hamiltonian and energy minimizers under conservative constraints, and as perturbations of simpler objects like non-rotating Lane-Emden stars or relative equilibria for point masses (e.g., [23, Section 3.2]). This characterization is essential for showing that the support of the energy constrained minimizer $\rho$ is contained within a ball of a certain radius, making $\rho$ a local minimizer with respect to the topology induced by Wasserstein $L^{\infty}$ (denoted by $W^{\infty}$ hereafter) distance. As discussed in Theorem 2.25 (see also [31, Theorem 2.1]), this ensures the existence of $\nabla P(\rho)$ in $\mathbb{R}^{3}$ and validates (EP’) in the whole space. However, McCann does not thoroughly address the existence of $\nabla P(\rho)$. While McCann provides valuable insights into the $W^{\infty}$ distance, further exploration is warranted—particularly why the local minimizer in this topology has finite energy, enabling the existence of the variational derivatives. These are what we aim to refine in this paper.
To be precise, this paper makes the following contributions:
  1. (a)
    Gradient Existence and Equation Derivation: We establish the existence of $\nabla P(\rho)$ for local minimizer $\rho$, which enables a rigorous transition from the Euler-Lagrange equation to the Euler-Poisson equation. The result is formalized in Theorem 2.25 and proved in Section 3.
  2. (b)
    Existence of $L^{\infty}$ Functions: We prove Lemma 4.7 (vi), showing any neighborhood in the topology induced by $W^{\infty}$ distance contains $L^{\infty}$ functions.
  3. (c)
    Energy Finiteness and Comparison of Topologies: We furthermore establish the finiteness of energy and the variational derivatives for local minimizers in Lemma 5.7. We also discuss results in the alternative topology inherited from the topological vector space, where we show finite‑energy local minimizers do not exist (Proposition 5.15), yet weak local minimizers with infinite energy can be found (Proposition 5.23).
Part of the results presented in this paper are based on the author’s Master’s thesis [12]. Beyond their intrinsic mathematical interest, these results serve as preliminary findings and will be used in the author’s companion paper [13] on star–planet systems, which also originates from the same thesis. In [13], they serve as an alternative to direct citations of McCann’s conclusions [31], providing a more precise and rigorous foundation for the arguments therein.
The structure of this paper is outlined as follows: in Section 2, we review McCann’s constructions and results of binary stars [31], with some nontrivial modifications. In Section 3, we discuss the necessity of those nontrivial modifications, provide nontrivial supplementary details regarding the properties of local minimizers under the topology induced by $W^{\infty}$ distance, particularly focusing on contribution (a). In Section 4, we address contribution (b), establishing the existence of $L^{\infty}$ functions in any neighborhood. In Section 5, we accomplish contribution (c): we apply results in Section 4 to deduce the finiteness of energy for local minimizers and compare it with the results in alternative topology.
2 McCann’s Construction and Results of 2-Body Systems
In this section we review McCann’s construction and results of 2-Body systems [31]. In the first subsection, we introduce some basic settings, notations, assumptions and problem settings such as variational formulation. In the second subsection, we focus on explaining why the general energy form $E(\rho,v)$ can be replaced by the energy with uniform rotation $E_{J}(\rho)$. In the third subsection, we present some of McCann’s nice conclusions.
2.1 Notations and problem setting
We introduce some notations and problem settings based on McCann’s one [31] and Jang and Seok’s [23].
Definition 2.1 (Notations).
We give the following definitions:
  1. (i)
    k{k} dimensional open ball: $B_{R}(a):=\left\{x\in\mathbb{R}^{k}\mid|x-a\mid<R\right\}$ for given $a\in\mathbb{R}^{k},R\in\mathbb{R}$.
  2. (ii)
    The projection operator of $x$ to the $x_{1}x_{2}$ plane: $P_{12}(x)=P_{12}\left(x_{1},x_{2},x_{3}\right):=\left(x_{1},x_{2},0\right)$.
  3. (iii)
    A bilinear form $\langle\cdot,\cdot\rangle_{2}:\mathbb{R}^{3}\times\mathbb{R}^{3}\rightarrow\mathbb{R}:\forall x,y\in\mathbb{R}^{3},\langle x,y\rangle_{2}=P_{12}(x)\cdot P_{12}(y)=x_{1}y_{1}+x_{2}y_{2}$.
  4. (iv)
    $r(x):=\left(\langle x,x\rangle_{2}\right)^{\frac{1}{2}}=x_{1}^{2}+x_{2}^{2}$.
  5. (v)
    Rotation map: R_θ:=(cosθ-sinθ0sinθcosθ0001)
  6. (vi)
    Nonnegative (Positive) part $\lambda]_{+}:=\max\{\lambda,0\}$.
  7. (vii)
    $W^{\infty}$ metric: Wasserstein $L^{\infty}$ metric.
  8. (viii)
    Given $\delta>0$ the $\delta$-neighbourhood of $\Omega\subset\mathbb{R}^{3}$ to be the set Ω+B_δ(0):=⋃_y ∈Ω{x ∈R^3∣— x-y —¡δ}
  9. (ix)
    $\mu$: Lebesgue measure.
The state of a fluid may be represented by its mass density $\rho(x)\geq 0$ and velocity vector field $v(x)$ on $\mathbb{R}^{3}$. The fluid interacts with itself through Newtonian gravity hence we need to consider gravitational interaction energy, which will be given later. Moreover, to define internal energy, we follow Auchmuty and Beals’s assumptions [5] and first consider a general form of the pressure $P(\rho)$ as the following:
  1. (F1)
    $P:[0,\infty)\rightarrow[0,\infty)$ continuous and strictly increasing;
  2. (F2)
    $\lim\limits_{\rho\rightarrow 0}P(\rho)\rho^{-\frac{4}{3}}=0$
  3. (F3)
    $\lim\limits_{\rho\rightarrow\infty}P(\rho)\rho^{-\frac{4}{3}}=\infty$.
With these assumptions, we also define $A(\rho)$ as the following:
$A(\rho):=\int_{1}^{\infty}P\left(\frac{\rho}{v}\right)dv=\rho\int_{0}^{\rho}P(\tau)\tau^{-2}d\tau$ (1)
Due to (F2) we know $A(\rho)$ is well-defined. Easy to see $A(\rho)$ is strictly increasing. It turns out that $A(\rho)$ is also convex (see Lemma 2.6), and it is related to the internal energy $U(\rho)$ given in (7) below. Moreover, if polytropic equations of state $P(\rho)=K\rho^{\gamma}$ holds, where the parameter $\gamma>\frac{4}{3}$, then easy to check $P$ satisfies (F1) (F2) (F3), and
$A(\rho)=\frac{K}{\gamma-1}\rho^{\gamma}$ (2)
Remark 2.2.
When showing the existence of non-rotating single star, (F3) can be replaced by the assumption
$\lim\limits_{\rho\rightarrow\infty}\inf P(\rho)\rho^{-\frac{4}{3}}>K$ (F3’)
where $K>0$ should be large enough to prevent gravitational collapse, and thereby the Chandrasekhar mass for the model is assumed greater than $m=1$. Then we can obtain minimizers with total mass not larger than 1 via variational method. See [31, Section 2] or [5, Section 1,6,8 and Appendix].
Remark 2.3.
By L’Hôpital’s rule [35, 5.13 Theorem], and the integration expression of $A(s)=s\int_{0}^{s}P(t)t^{-2}dt$ in (1), we can prove that AA also satisfies (F2) and (F3) (or (F3’) if $P$ satisfies (F3’) instead of (F3), though the constant $K$ can be different).
By (1), we can obtain $A^{\prime}(s)$ satisfies
$A^{\prime}(s)=\left\{\begin{array}[]{r}\int_{0}^{s}P(t)t^{-2}dt+\frac{P(s)}{s},s>0\\ 0,s=0\end{array}\right$. (3)
Notice $A^{\prime}(s)$ is continuous and $A^{\prime}(s)s-A(s)=P(s)$.
Remark 2.4.
By (3), we can also prove that $A^{\prime}$ satisfies: $\lim\limits_{\rho\rightarrow 0}A^{\prime}(\rho)\rho^{-\frac{1}{3}}=0$ and $\lim\limits_{\rho\rightarrow\infty}A^{\prime}(\rho)\rho^{-\frac{1}{3}}=\infty$.
Remark 2.5.
We claim the assumptions above can imply $A$ is convex and $A^{\prime}$ is strictly increasing. One formal way to see it, as Auchmuty and Beals suggest [5], is by taking the derivative again. Due to (F1), we know $P^{\prime}(s)$ exists almost everywhere and $P^{\prime}$ is locally integrable, (one can refer to [34, Lebesgue’s Theorem and Corollary 4 below]). Since $A^{\prime}(s)s-A(s)=P(s)$, it is conceivable $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}\geq 0$ a.e., which might imply $A^{\prime}(s)$ increases, and then $A(s)$ is convex [35, Chapter 5, Exercise 14]. However, it is not really sufficient. One counterexample can be $f(s)=-\Theta(s)$, where $\Theta(s)$ is Cantor-Lebesgue function, which means $f^{\prime}(s)=0$ a.e. while $f$ decreases and is not constant. Therefore, we need a more detailed analysis to prove rigorously that $A^{\prime}$ is strictly increasing.
Lemma 2.6 (Strictly Increasing of $A^{\prime}$).
If $P(\rho)$ satisfies assumptions (F1) (F2) (F3), then $A^{\prime}$ is continuous and strictly increasing, where $A$ is given in (1). Moreover, $A$ is convex.
Proof.
Due to (3), we can see $A^{\prime}$ is continuous and $A^{\prime}(s)>A^{\prime}(0)=0$ when $s>0$. Given $s>0$ and $h>0$, by (3) we have $A^{\prime}(s+h)-A^{\prime}(s)=\int_{s}^{s+h}P(t)t^{-2}dt+\frac{P(s+h)}{s+h}-\frac{P(s)}{s}$. Since $P$ is monotonic (strictly increasing), by second mean value theorem for definite integrals [20], we know $\exists\xi\in(0,h)$, $\int_{s}^{s+h}P(t)t^{-2}dt=P(s)\int_{s}^{s+\xi}t^{-2}dt+P(s+h)\int_{s+\xi}^{s+h}t^{-2}dt=P(s)\left(\frac{1}{s}-\frac{1}{s+\xi}\right)+P(s+h)\left(\frac{1}{s+\xi}-\frac{1}{s+h}\right)$. Then $A^{\prime}(s+h)-A^{\prime}(s)=\frac{P(s+h)-P(s)}{s+\xi}>0$, which means $A^{\prime}$ is strictly increasing on $0,\infty)$. The convexity of $A$ can refer to [35, Chapter 5, Exercise 14]
Remark 2.7.
Thanks to Lemma 2.6, the inverse function of $A^{\prime}$, denoted by $\left(A^{\prime}\right)^{-1}$ or $\phi$, is well defined on $0,\infty)$. Moreover, $\phi=\left(A^{\prime}\right)^{-1}$ is continuous since $A^{\prime}$ is continuous.
Another idea to show the existence of $\phi$, at least local existence, is using inverse function theorem. Assume $P(s)$ is continuously differentiable, we can know when $s>0,A^{\prime\prime}(s)$ exists and $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}$. However, one could run into trouble when using inverse function theorem if $P^{\prime}(s)=0$ (notice $P$ strictly increases does not mean $P^{\prime}(s)>0$, for example $f(x)=x^{3}$ at point $x=0$). But since the differentiability of $P$ is one idea to show the existence of $\nabla P(\rho)$, which is one term in the Euler-Poisson equation (EP) (or (EP’)), we can make an additional assumption of $P(\rho)$:
  1. (F4)
    $P(\rho)$ is continuously differentiable on $0,\infty)$, and $P^{\prime}(\rho)>0$ if $\rho>0$.
Remark 2.8.
The motivation of assuming (F4) is the following: later the minimizer $\sigma$ satisfies $\sigma=\phi\circ W_{\sigma}$ where $\sigma>0$ ($W_{\sigma}$ will be specified later), and then the pressure $P(\sigma)=P\circ\phi\circ W_{\sigma}$ in $\{\sigma>0\}$. By (F4), we know $A^{\prime\prime}(\rho)$ exists and $A^{\prime\prime}(\rho)=\frac{P^{\prime}(\rho)}{\rho}\neq 0$ is continuous if $\rho>0$, which implies $\phi=\left(A^{\prime}\right)^{-1}\in C^{1}((0,\infty))$ (one can refer to [38, Theorem 10.4.2]). It turns out this can help to show the differentiability of $\sigma$ and then of $P(\sigma)$ where $\sigma>0$, as we can see in Theorem 2.25 below.
Since actually what we want is the differentiability of $P(\sigma)$ rather than of $\sigma$ to show (EP’), another alternative assumption, which is more general but also somewhat technical, is:
  1. (F4’)
    $P(\rho)$ is continuously differentiable on $0,\infty)$. If $\rho>0,P(\rho)$ has non-vanishing (first order or higher order) derivative at $\rho$. That is, $\exists n\geq 1$, such that $P^{(n)}(\rho)$ exists and is not 0.
The advantages and applications of this more general assumption will be demonstrated later, such as in Lemma 3.1 and Theorem 2.25 (vii).
In the following sections, we assume $P(\rho)$ satisfies (F1)(F2)(F3) unless otherwise specified. We will mention (F4) or (F4’) or other assumptions when we want to use it. As a special case, if we assume the polytropic equations of state with index $\gamma>\frac{4}{3}$ holds true, one can check the assumptions (F1)(F2)(F3)(F4) or (F1)(F2)(F3)(F4’) above are satisfied automatically.
We consider a “potential admissible class” for $\rho$ and $v$ as the following:
$R\left(\mathbb{R}^{3}\right)$ $:=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int\rho=1\right\}$ (4)
[…]
Then given $\rho$ and $v$ in such sets, McCann [31] consider energy $E(\rho,v)$ consisting of three terms:
$E(\rho,v):=U(\rho)-\frac{G(\rho,\rho)}{2}+T(\rho,v)$ (6)
$U(\rho):=\int A(\rho(x))d^{3}x$ (7)
$G(\sigma,\rho):=\int V_{\sigma}\rho=\iint\frac{\sigma(x)\rho(y)}{|x-y|}d^{3}yd^{3}x$ (8)
$T(\rho,v):=\frac{1}{2}\int|v|^{2}\rho$ (9)
Here $A(\rho)$ is a convex function given in (1), and $U(\rho)$ is internal energy. $G$ is viewed as a symmetric bilinear form since this observation can help us to compute the variational derivative in Lemma 5.3 in subsection 5.1. Then the gravitational potential energy is defined as $G(\rho,\rho)$.
We can choose units so that the total mass of fluid is one and the gravitational constant $G=1$, then $V_{\rho}$ represents the gravitational potential of the mass density $\rho(x)$
$V_{\rho}(x):=\int\frac{\rho(y)}{|y-x|}d^{3}y$ (10)
Hence $G(\rho,\rho)$ is gravitational interaction energy.
We can also see $T(\rho,v)$ is kinetic energy.
We can choose a frame of reference in which the center of mass
$\bar{x}(\rho):=\frac{\int x\rho(x)d^{3}x}{\int\rho(x)d^{3}x}$ (11)
is at rest. We are interested in finding minimum energy configurations subject to constraints of fixed mass ratio and angular momentum $\boldsymbol{J}$ with respect to the center of mass $\bar{x}(\rho)$. The fluid angular momentum is given by $\boldsymbol{J}(\rho,v)$:
$\boldsymbol{J}(\rho,v):=\int(x-\bar{x}(\rho))\times v\rho(x)d^{3}x$ (12)
We denote by $J_{z}$ the z-component of $\boldsymbol{J}$, that is, $J_{z}(\rho,v):=\hat{e}_{z}\cdot\boldsymbol{J}(\rho,v)$, where $\hat{e}_{z}=(0,0,1)^{T}$. For simplicity of notation, we will sometimes use $J$ to represent $J_{z}$ and call $J$ the angular momentum of the system when no confusion arises.
Since the z-component of the angular momentum is specified, the moment of inertia $I(\rho)$ of $\rho$ in the direction of $\hat{e}_{z}$ will be relevant. That is, we define $I(\rho)$ to be moment of inertia of $\rho$ in the direction of $\hat{e}_{z}$:
$I(\rho):=\int\rho r^{2}(x-\bar{x}(\rho))$ (13)
Remark 2.9.
When $\rho$ has positive mass, we have $I(\rho)>0$. If not, $I(\rho)=\int\rho(x)r^{2}(x-\bar{x}(\rho))=0$ implies $\rho(x)r^{2}(x-\bar{x}(\rho))=0$ almost everywhere in $\mathbb{R}^{3}$. Since $\mu\left(\left\{x\in\mathbb{R}^{3}\mid r^{2}(x-\bar{x}(\rho))=0\right\}\right)=0$, we have $\rho(x)=0$ almost everywhere in $\mathbb{R}^{3}$, which contradicts the fact that $\rho$ has positive mass.
A simple case is the non-rotating problem, i.e. $\boldsymbol{J}=0$, with energy $E_{0}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}$. Properties of non-rotating minimizers are introduced in [14, 31, 5, 29].
To ensure that the definition of $E(\rho,v)$ (or $E_{J}(\rho)$ given later) does not result in a situation of $\infty$-$\infty$, we hope the gravitational potential energy $G(\rho,\rho)$ to be finite. Thanks to our setup, this is indeed the case, which leads to the following proposition:
Proposition 2.10 (Finite Gravitational Interaction Energy).
Suppose $\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, with $\left\|\rho\right\|_{L^{\frac{4}{3}}{(\mathbb{R}^{3})}}+\left\|\rho\right\|_{L^{1}{(\mathbb{R}^{3})}}<\lambda$. Then there is a constant $C(\lambda)$ which depends only on $\lambda$, such that $G(\rho,\rho)<C(\lambda)$.
Proof.
Since $\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, by Hardy-Littlewood-Sobolev Inequality Proposition A.7 we have $V_{\rho}\in L^{12}\left(\mathbb{R}^{3}\right)$, with $\left\|V_{\rho}\right\|_{L^{12}\left(\mathbb{R}^{3}\right)}<C_{1}(\lambda)$ for some constant $C_{1}(\lambda)>0$. By Interpolation Inequality Proposition A.6 we have $\rho\in L^{\frac{12}{11}}\left(\mathbb{R}^{3}\right)$, with $\|\rho\|_{L^{\frac{12}{11}\left(\mathbb{R}^{3}\right)}}<C_{2}(\lambda)$ for some constant $C_{2}(\lambda)>0$. Finally, by Hölder’s inequality we have
$G(\rho,\rho)=\int\rho V_{\rho}\leq\|\rho\|_{L^{\frac{12}{11}\left(\mathbb{R}^{3}\right)}}\left\|V_{\rho}\right\|_{L^{12}\left(\mathbb{R}^{3}\right)}<C_{1}(\lambda)\cdot C_{2}(\lambda)\coloneq C(\lambda)$
Remark 2.11.
With similar arguments, we can show a more general result: suppose $\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$ and $\sigma\in L^{\frac{12}{11}}\left(\mathbb{R}^{3}\right)$, then $G(\rho,\sigma)<\infty$.
Remark 2.12.
From Proposition 2.10, we see one reason why we require $\rho$ to belong not only to $L^{1}$ but also to $L^{\frac{4}{3}}$. Furthermore, the rationale behind selecting the specific exponent $\frac{4}{3}$ can also refers to McCann’s construction, which is a natural assumption by interpolation inequality Proposition A.6 if we hope $\inf\limits_{\rho\in{R}\left(\mathbb{R}^{3}\right)}E_{0}(m\rho)$ to be finite, see [5, Section 1,6 and 8], and [31, Section 2].
Consider the non-rotating minimizer $\sigma_{m}$ of $E_{0}(\rho)$ among configurations of mass $m\leq 1$, the corresponding minimum energy is finite due to the remark above. For the sake of convenience, we denote them by
$e_{0}(m):=E_{0}\left(\sigma_{m}\right)=\inf_{\rho\in{R}\left(\mathbb{R}^{3}\right)}E_{0}(m\rho).$ (14)
Sometimes $e_{0}(1)$ is denoted by $e_{0}$.
Intuitively we hope the minimizer should have a center of mass and finite moments of inertia, and one can see equation (EP’) will be satisfied automatically if $V_{\rho}$ has finite gradient and $x$ is the exterior point of the support of $\rho$ (denoted by spt $\rho$) as demonstrated in the proof of Theorem 2.25 (vii) in Section 3. Here support of $\rho$ is the smallest closed set carrying the full mass of $\rho$.
Therefore, we also introduce a centered subset $R_{a}\left(\mathbb{R}^{3}\right)$ of $R\left(\mathbb{R}^{3}\right)$:
${R}_{a}\left(\mathbb{R}^{3}\right):=\left\{\rho\in{R}\left(\mathbb{R}^{3}\right)\mid\bar{x}(\rho)=a;\text{ spt }\rho\text{ is bounded.}\right\}$ (15)
In general, it is easier to consider the case that the density is centered at 0, which is in ${R}_{0}\left(\mathbb{R}^{3}\right)$.
In subsection 2.2 one can see when the topology satisfies certain conditions, if the $z$-component $J=J_{z}$ of the angular momentum $\boldsymbol{J}$ is fixed, the problem of minimizing $E(\rho,v)$ locally in $R_{0}(\mathbb{R}^{3})\times V(\mathbb{R}^{3})$ is equivalent to a minimization problem of the energy $E_{J}(\rho)$ locally in $R_{0}(\mathbb{R}^{3})$. Here $E_{J}(\rho)$ corresponds to uniform rotation with the angular momentum $\boldsymbol{J}=J\hat{e}_{z}=(0,0,J)^{T}$ specified a priori, and $E_{J}(\rho)$ is defined as the following:
$E_{J}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}+T_{J}(\rho)$ (16)
$T_{J}(\rho):=\frac{J^{2}}{2I(\rho)}$
Due to this uniform rotation observation, in the presence of 2-body systems, although the systems rotate in universe, after fixing a time, we can still assume that they fall within two disjoint regions $\Omega_{m}$ and $\Omega_{1-m}$, widely separated relative to $\frac{J^{2}}{\mu^{2}}$, where $\mu=m(1-m)$ is their reduced mass. Note that the same symbol $\mu$ is also used as for Lebesgue measure, provided no ambiguity arises. For the planet’s mass $m\in(0,1)$ and the star’s mass (1m)(1-m), we consider $E_{J}(\rho)$ is minimized subject to the constraint
$W_{m,J}\coloneq\left\{\rho(m)=\rho_{m}+\rho_{1-m}\in R\left(\mathbb{R}^{3}\right)\mid\int\rho_{m}=m,\text{spt }\rho_{m}\subset\Omega_{m},\text{spt }\rho_{1-m}\subset\Omega_{1-m}\right\}$ (17)
where $\Omega_{m}$ and $\Omega_{1-m}$ are subsets of $\mathbb{R}^{3}$, which are given in the following. The subscripts $m$ and $J$ in $W_{m,J}$ indicate $\Omega_{m}$ and $\Omega_{1-m}$ are related to $m$ and $J$.
Fix two points $y_{m}$ and $y_{1-m}$ in $\mathbb{R}^{3}$ from the plane $z=0$, which are separated by $\eta=\frac{J^{2}}{\mu^{2}}$, i.e. $\eta=\left|y_{m}-y_{1-m}\right|$. The $\Omega_{m}$ and $\Omega_{1-m}$ are defined as closed balls in $\mathbb{R}^{3}$ centered at $y_{m}$ and $y_{1-m}$, whose size and separation scale with $\eta$ as the following:
$\Omega_{m}$ $=\left\{x\in\mathbb{R}^{3}\mid|x-y_{m}|\,\leq\frac{\eta}{4}\right\}$ (18)
$\Omega_{1-m}$ $=\left\{x\in\mathbb{R}^{3}\mid|x-y_{1-m}|\,\leq\frac{\eta}{4}\right\}$
The distance separating $\Omega_{m}$ and $\Omega_{1-m}$, and the diameter of their union is given by:
${dist}\left(\Omega_{m},\Omega_{1-m}\right)$ $=\frac{\eta}{2}$ (19)
${diam}\left(\Omega_{m},\Omega_{1-m}\right)$ $=\frac{3\eta}{2}$ (20)
Remark 2.13.
The reason we set the separation $\eta=\frac{J^{2}}{\mu^{2}}$ in the definitions above is inspired by the Kepler problem. Given two point masses $m$ and 1m1-m, rotating with angular momentum $J>0$ about their fixed center of mass. Assume their separation is $d$, the gravitational energy plus kinetic energy is $\frac{\mu}{d}+\frac{J^{2}}{2\mu d^{2}}$, which reaches its minimum at separation $d=\eta$.
Remark 2.14.
We hope that the radii of $\Omega_{m}$ and $\Omega_{1-m}$ increase as $\eta$ increases is to guarantee the supports of constrained minimizers (stars or planets) will fall in the interior of $\Omega_{m}\cup\Omega_{1-m}$. This is because McCann shows the size of stars or planets will not expand too much as $\eta$ increases. As discussed in [31, Section 6], it lays the groundwork for discussion on the constrained minimizers being local minimizers (here and in what follows, unless otherwise specified, “local” will always refer to “local under the topology induced by the Wasserstein $L^{\infty}$ distance”). Therefore, the minimizers are solutions to (EP’) due to Theorem 2.25.
When the distance between two objects is large, the gravitational interaction between them becomes very weak. As a result, the stars are not significantly deflected, and their supports remain limited. Therefore, they can be approximated as point masses, which simplifies the analysis.
When $\eta$ is large (which occurs when $\mu$ is small or $J$ is large), since the gravitational interaction between two bodies seems small, the stars are not significantly deflected, and their sizes are negligible compared to the distance between them. Hence, one can expect a stable, slowly rotating equilibrium to exist in which fluid components with masses $m$ and 1m1-m lie near $y_{m}$ and $y_{1-m}$. It is indeed true as McCann discussed in [31, Section 6]. For prescribed angular momentum, we also know the energy $E(\rho,v)$ is bounded from below on ${R}_{0}\left(\mathbb{R}^{3}\right)$ by the non-rotating energy $e_{0}$. However, as in Morgan [32], McCann [31, Example 3.6] demonstrates that this bound — although approached — will not be attained, see also [30]. Therefore, instead of searching for a global energy minimizer in binary-star system, one is forced to settle for local minimizers of $E(\rho,v)$ in an appropriate topology.
In fact, the velocity fields $v$ may be topologized in any way which makes ${V}\left(\mathbb{R}^{3}\right)$ a topological vector space, since the velocity distribution minimizing the kinetic energy is always given by uniform rotation [31, Section 3] [23, Lemma 2.2.2] as we will describe in subsection 2.2. However, the choice of topology for ${R}_{0}\left(\mathbb{R}^{3}\right)$ is quite delicate: McCann pointed out local energy minimizer will not exist if the topology of $R_{0}(\mathbb{R}^{3})$ is inherited from a topological vector space ([31, Remark 3.7]), i.e. ${R}_{0}\left(\mathbb{R}^{3}\right)$ is a subset of a topological vector space (see [33, Chapter 1, Section 16]). While we partially acknowledge this perspective, we argue that the focus should be on $R(\mathbb{R}^{3})$ rather than $R_{0}(\mathbb{R}^{3})$. For further details, one can refer to Proposition 5.15 and Remark 5.19 in subsection 5.2. We will further compare this case with the case that the topology induced by the Wasserstein $L^{\infty}$ distance, to highlight the advantages of considering topology induced by the Wasserstein $L^{\infty}$ distance in such problems.
Since the star and planet are separated, it is convenient to describe the relations between the total moment of inertia $I\left(\rho_{m}+\rho_{1-m}\right)$ and $I\left(\rho_{m}\right),I\left(\rho_{1-m}\right)$.
Lemma 2.15 (Expansion of Moment of Inertia).
Let $\rho\geq 0$, $\sigma\geq 0$ be the density functions in $\mathbb{R}^{3}$ with mass $\int\rho=m_{1}<\infty,\int\sigma=m_{2}<\infty$. $\bar{x}(\rho)$ and $\bar{x}(\sigma)$ denote the centers of mass, $I(\rho)$ and $I(\sigma)$ denote the moments of inertia, function $r$ is given in (2.1).
  1. (1)
    If $m_{1}+m_{2}=0$, then $I(\rho+\sigma)=0$.
  2. (2)
    If $m_{1}+m_{2}>0$, then we have the moment of inertia of $\rho+\sigma$ satisfies
    $I(\rho+\sigma)=I(\rho)+I(\sigma)+\frac{m_{1}m_{2}}{m_{1}+m_{2}}r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))$ (21)
Proof.
The (1) case is obvious, since $m_{1}+m_{2}=0$ implies $\rho+\sigma$ is 0 almost everywhere. For (2) case, by definition $\left(m_{1}+m_{2}\right)\bar{x}(\rho+\sigma)=\int(\rho+\sigma)x=m_{1}\bar{x}(\rho)+m_{2}\bar{x}(\sigma)$, then
$I(\rho+\sigma)$ $=\int(\rho+\sigma)r^{2}(x-\bar{x}(\rho+\sigma))$
$=\int\rho r^{2}(x-\bar{x}(\rho)+\bar{x}(\rho)-\bar{x}(\rho+\sigma))+\int\sigma r^{2}(x-\bar{x}(\sigma)+\bar{x}(\sigma)-\bar{x}(\rho+\sigma))$
Consider the first term on the right above, we have
$\int\rho r^{2}(x-\bar{x}(\rho)+\bar{x}(\rho)-\bar{x}(\rho+\sigma))=$ $\int\rho r^{2}(x-\bar{x}(\rho))+\int\rho r^{2}(\bar{x}(\rho)-\bar{x}(\rho+\sigma))$
$+2\int\rho\langle x-\bar{x}(\rho),\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}$
Notice by definition of center of mass, we have $\int\rho\langle x,\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}=m_{1}\langle\bar{x}(\rho),\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}=\int\rho\langle\bar{x}(\rho),\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}$, thus $2\int\rho\langle x-\bar{x}(\rho),\bar{x}(\rho)-\bar{x}(\rho+\sigma)\rangle_{2}=0$, therefore,
$\int\rho r^{2}(x-\bar{x}(\rho)+\bar{x}(\rho)-\bar{x}(\rho+\sigma))$ $=I(\rho)+m_{1}r^{2}\left(\frac{m_{2}}{m_{1}+m_{2}}(\bar{x}(\rho)-\bar{x}(\sigma))\right)$
$=I(\rho)+\frac{m_{1}\cdot m_{2}^{2}}{\left(m_{1}+m_{2}\right)^{2}}r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))$
Similarly
$\int\sigma r^{2}(x-\bar{x}(\sigma)+\bar{x}(\sigma)-\bar{x}(\rho+\sigma))=I(\sigma)+\frac{m_{2}\cdot m_{1}^{2}}{\left(m_{1}+m_{2}\right)^{2}}r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))$
Thus
$I(\rho+\sigma)$ $=I(\rho)+\frac{m_{1}\cdot m_{2}^{2}}{\left(m_{1}+m_{2}\right)^{2}}r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))+I(\sigma)+\frac{m_{2}\cdot m_{1}^{2}}{\left(m_{1}+m_{2}\right)^{2}}r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))$
$=I(\rho)+I(\sigma)+\frac{m_{1}m_{2}}{m_{1}+m_{2}}r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))$
This is what we want. ∎
Remark 2.16.
Physically it’s natural to assume those mass, center of mass and moment of inertia finite. However, one may notice as long as the mass $m_{1}$ and $m_{2}$ are finite and one of the centers of mass is finite, the result can still hold even though the values may be infinite.
2.2 Uniform rotation minimizes kinetic energy
This subsection reviews several results of McCann [31], Jang and Seok [23]. First, we introduce that the problem of minimizing $E(\rho,v)$ is equivalent to a minimization problem of the energy $E_{J}(\rho)$ with respect to uniform rotation, that is, the fluid rotates uniformly about its center of mass. Then we show that with uniform rotation setting, the conservation of mass $\partial_{t}\rho+\nabla(\rho v)=0$ automatically holds true and then the Euler-Poisson equations (EP) can be simplified to (EP’) as Jang and Seok describes [23].
When we define the total energy $E(\rho,v)$, we can notice that the velocity fields $v$ only appears in the kinetic energy $T(\rho,v)$. Therefore, we can first can consider the velocity minimizer of $T(\rho,v)$ with a given density $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$. McCann points out that $T(\rho,v)$ is uniquely minimized by a uniform rotation:
Proposition 2.17 (Uniform Rotation around Center of Mass [28] [31, Proposition 3.1]).
Fix a fluid density $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$ and $J\geq 0$. Among all velocities $v\in{V}\left(\mathbb{R}^{3}\right)$ for which $T(\rho,v)<\infty$ and satisfying the constraint ${J}_{z}(\rho,v)=J$, the kinetic energy $T(\rho,v)$ is uniquely minimized by a uniform rotation $v(x):=\omega\hat{e}_{z}\times x$ with angular velocity $\omega:=\frac{J}{I(\rho)}$.
Proof.
Given $\rho$ in ${R}_{0}\left(\mathbb{R}^{3}\right)$ fixed, we define the Hilbert space $H\coloneq L^{2}\left({\mathbb{R}^{3},d\rho(x)}\right)\subset V\left(\mathbb{R}^{3}\right)$, where $d\rho(x)=\rho(x)d^{3}x$, with inner product $v,w>_{H}\coloneq\int(v\cdot w)\rho$, in particular $v,v>_{H}=2T(\rho,v)$. One can check: (1) $\hat{e_{z}}\times x\in H$; (2) if $v\in{V}\left(\mathbb{R}^{3}\right)$ satisfies $T(\rho,v)<\infty$ and the constraint ${J}_{z}(\rho,v)=J$, then it is in the space $G=\left\{v\in H\middle|<v,~\hat{e_{z}}\times x>_{H}=J\right\}$; (3) if $v$ is in $G$, then $v$ can be expressed as $v=v_{g}+v_{o}$, where $v_{g}=\frac{J}{I(\rho)}\left(\hat{e_{z}}\times x\right)=\omega\left(\hat{e_{z}}\times x\right)$ and $v_{o}$ is orthogonal to $\hat{e_{z}}\times x$ in $H$, that is $v_{0},\hat{e_{z}}\times x>_{H}=0$. Therefore,
$<v,v>_{H}$ $=<v_{g},v_{g}>_{H}+<v_{0},v_{0}>_{H}$
$\geq<v_{g},v_{g}>_{H}$
$=2T\left(\rho,v_{g}\right)$
For the detailed proof, one can refer to [31, Proposition 3.1]. ∎
If $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$ rotates with velocity $v(x):=\frac{J}{I(\rho)}\hat{e}_{z}\times x$, then it is easy to compute that its kinetic energy $T(\rho,v)$ is given by
$T_{J}(\rho):=\frac{J^{2}}{2I(\rho)}$ (22)
Thanks to Remark 2.9, we know $T_{J}(\rho)=\frac{J^{2}}{2I(\rho)}$ is well-defined and finite if the mass of $\rho$ is not zero.
We have the following corollary:
Corollary 2.18 (Velocity-free Reformulation [31, Secion 3]).
Let $J\geq 0$ and $\rho\in{R}\left(\mathbb{R}^{3}\right)$, and define $\omega:=\frac{J}{I(\rho)}$. Let $V(\mathbb{R}^{3})$ be a topological vector space, and the topology on ${R}_{0}\left(\mathbb{R}^{3}\right)$ is such that the map taking $\rho$ to $I(\rho)$ is continuous. Then $(\rho,v)$ minimizes $E(\rho,v)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ subject to the constraint $J_{z}(\rho,v)=J$ if and only if $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)$ and $v(x)=\omega\hat{e}_{z}\times x$. Here $E_{J}(\rho)=U(\rho)-\frac{G(\rho,\rho)}{2}+T_{J}(\rho)$.
Remark 2.19.
By Lemma 4.7 (v), we know that ${R}_{0}\left(\mathbb{R}^{3}\right)$ is endowed with the topology induced by the Wasserstein $L^{\infty}$ distance, then it satisfies the conditions required in the above corollary, namely, that the mapping $\rho\mapsto I(\rho)$ is continuous.
Remark 2.20.
Corollary 2.18 implies even if $(\rho,v)$ minimizes $E(\rho,v)$ only locally on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ subject to the constraint $J_{z}(\rho,v)=J$, we still have $v(x)=\omega\hat{e}_{z}\times x$, which, to some extent, strengthens the result of Proposition 2.17. Actually it comes from McCann’s result [31, Corollary 3.2], where the minimum of $T(\rho,\widehat{v}(t))$ or $E(\rho,\widehat{v}(t))$ along the curve $\widehat{v}(t)=(1-t)\tilde{v}(x)+t\omega\hat{e}_{z}\times x$ is considered.
Remark 2.21.
In the arguments above, we only consider minimization on ${R}_{0}\left(\mathbb{R}^{3}\right)$ or ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$. By translation invariant of energy we can see the results also holds over ${R}_{a}\left(\mathbb{R}^{3}\right)$ or ${R}_{a}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ for any $a\in\mathbb{R}^{3}$, and furthermore, it can actually be extended to minimization over ${R}\left(\mathbb{R}^{3}\right)$ or ${R}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$. However, since the solutions to (EP’) considered later are in $W_{m,J}$, in particular, they have compact support, hence the above result suffices for our purposes.
In summary, the local minimization in ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ of $E(\rho,v)$ is equivalent to the local minimization over ${R}_{0}\left(\mathbb{R}^{3}\right)$ of $E_{J}(\rho)$. Note the internal energy and gravitational potential energy are rotation‑invariant. Under the constraint $J_{z}(\rho,v)=J$, when we know $\widetilde{\rho}$ is a local minimizer over ${R}_{0}\left(\mathbb{R}^{3}\right)$ of $E_{J}(\rho)$, we expect—on physical grounds—that the celestial bodies should be rotating rather than stationary. Therefore, we can further define
$({\rho}(t,x),v(t,x)):=\left(\widetilde{\rho}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$
where $\omega=\frac{J}{I\left(\rho\right)}$, then $({\rho}(t,x),v(t,x))$ gives a uniform rotating binary star system (or star-planet system) such that $E(\rho,v)$ is local minimum.
Now we have understood one motivations of defining $E_{J}$. Actually, $E_{J}$ can be derived through another approach, inspired by Jang and Seok [23].
To find rotating $N$-body solutions to the Euler-Poisson system, we may assume $\rho(t,y)$, $v(t,y)$, and $V(t,y)$ have the following forms:
$\rho(t,y)=\widetilde{\rho}\left(R_{-\omega t}y\right),v(t,y)=\omega\left(-y_{2},y_{1},0\right)^{T},V(t,y)=V_{\widetilde{\rho}}\left(R_{-\omega t}y\right)$
where $\omega>0$ is angular velocity, $\widetilde{\rho}$ is a nonnegative density function with compact support. Then we can see the Poisson equation is satisfied automatically. For the conservation of mass, we compute:
$\partial_{t}\rho$ $=\partial_{y_{1}}\widetilde{\rho}\cdot\left(-y_{1}\sin\omega t\omega+y_{2}\cos\omega t\omega\right)+\partial_{y_{2}}\widetilde{\rho}\cdot\left(-y_{1}\cos\omega t\omega-y_{2}\sin\omega t\omega\right)$
$\nabla(\rho v)$ $=-\omega y_{2}\left(\partial_{y_{1}}\widetilde{\rho}\cos\omega t-\partial_{y_{2}}\widetilde{\rho}\sin\omega t\right)+\omega y_{1}\left(\partial_{y_{2}}\widetilde{\rho}\cos\omega t+\partial_{y_{1}}\widetilde{\rho}\sin\omega t\right)$
Thus $\partial_{t}\rho+\nabla(\rho v)=0$, the conservation of mass holds true.
We recall $P(\rho)$ denotes the pressure where $\rho$ is a density function, and $P_{12}(y)$ denotes the projection operator, where $y\in\mathbb{R}^{3}$. For the Euler (momentum) equation, we know it becomes
$\omega^{2}\widetilde{\rho}\left(R_{-\omega t}y\right)P_{12}(y)+R_{-\omega t}^{T}\left(\nabla_{\widetilde{x}}P\left(\widetilde{\rho}\left(R_{-\omega t}y\right)\right)\right)-\widetilde{\rho}\left(R_{-\omega t}y\right)R_{-\omega t}^{T}\left(\nabla_{y}V_{p}\left(R_{-\omega t}y\right)\right)=0$ (23)
Where $\nabla_{y}$ means the gradient operator w.r.t. $y$, and $\nabla_{x}$ means the gradient operator with w.r.t. $\widetilde{x}$, where $\widetilde{x}=R_{-\omega t}y$ above.
By the change of variable $x=R_{-\omega t}y$, we have $P_{12}(x)=R_{-\omega t}P_{12}(y),P_{12}(y)=R_{-\omega t}^{-1}P_{12}(x)=R_{-\omega t}^{T}P_{12}(x)$, then we get the equation introduced in the Introduction:
$\omega^{2}\widetilde{\rho}(x)P_{12}(x)+\nabla P(\widetilde{\rho}(x))-\widetilde{\rho}(x)\left(\nabla V_{\widetilde{p}}(x)\right)=0$ (EP’)
Where $\nabla$ means the gradient operator w.r.t. $x$ by default. We call (EP’) reduced Euler-Poisson equation. One can also reverse this process to demonstrate that (EP’) can imply (EP) as well.
If we further assume $P^{\prime}$ and $\nabla\widetilde{\rho}$ exists, then (EP’) becomes
$\omega^{2}\widetilde{\rho}(x)P(x)+P^{\prime}(\widetilde{\rho}(x))\nabla(\widetilde{\rho}(x))-\widetilde{\rho}(x)\left(\nabla V_{\widetilde{p}}(x)\right)=0$ (24)
Since we consider $N$-body solutions, we denote
$\widetilde{\rho}=\sum_{i=1}^{N}\widetilde{\rho}_{i},\quad\widetilde{\rho}_{i}\geq 0$
such that $\text{spt}\left(\widetilde{\rho}_{i}\right)$ is connected for every $i\in 1,\ldots,N$ and mutually disjoint. Notice when $s>0$, $A^{\prime\prime}(s)$ exists and $A^{\prime\prime}(s)s=P^{\prime}(s)$ where $A$ is given in (1). Then in the region $\{\widetilde{\rho}>0\}$, dividing (24) by $\widetilde{\rho}$, one has
$\nabla\left(-\frac{1}{2}\omega^{2}|P_{12}(x)|^{2}+A^{\prime}\left(\widetilde{\rho}_{i}\right)-V_{\widetilde{p}}\right)=0,\forall x\in\left\{\widetilde{\rho}_{i}>0\right\}$
Therefore, we have
[…]
for some constants $C_{i},i=1,\ldots,N$.
[…]
Remark 2.22.
Someone might be a bit confused when they first see the integration definition of $A(\rho)$ in (1). But it turns out with this definition we can have $A^{\prime}(\rho)\rho-A(\rho)=P(\rho)$. If $P(\rho)$ is continuous differentiable, we can have $A^{\prime\prime}(\rho)\rho=P^{\prime}(\rho)$ where $\rho>0$, which has been used above to obtain (EL0’).
Remark 2.23.
One can notice in (23) we assume the density rotates uniformly around 0, which implicitly assumes that the center of mass is located at the origin. However, from a physical standpoint, it is more appropriate to consider the object rotating about its center of mass. Also, from a mathematical standpoint, thanks to the similar arguments in Proposition 2.17 and Corollary 2.18 above, and Theorem 2.25, we know that if $(\widetilde{\rho},v)$ is the minimizer, then the fluid should indeed rotate around its center of mass. Luckily, when we consider Euler-Lagrange equation (EL), although the corresponding minimizer does not necessarily have $(0,0,0)^{T}$ as its center of mass, we can apply similar arguments as above. Therefore, instead of (EL0’) sometimes it is more convenient for us to first consider the following kind of Euler-Lagrange equation:
[…]
Remark 2.24.
[…]
2.3 McCann’s results for 2-body systems
In this subsection, we review a slightly modified version of McCann’s results for Wasserstein $L^{\infty}$ ($W^{\infty}$) local energy minimizers and existence theorem of binary star solution (further discussion on modifications will be discussed in Section 3). We also make some remarks on them.
Although McCann assumes that pressure $P=P(\rho)$ satisfies (F1)(F2)(F3’), here we assume $P=P(\rho)$ satisfies (F1)(F2)(F3). Actually the arguments are mainly the same as we can see in Section 3.
Theorem 2.25 (Properties of $W^{\infty}$-Local Energy Minimizers [31, Theorem 2.1]).
Let $J>0$. If $(\rho,v)$ minimizes $E(\rho,v)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ subject to the constraint $J_{z}(\rho,v)=J$ then:
  1. (i)
    the z-axis is a principal axis of inertia for $\rho$, with a moment of inertia $I(\rho)$ which is maximal and non-degenerate;
  2. (ii)
    the rotation is uniform: $v(x):=\frac{J\hat{e}_{z}\times x}{I(\rho)}$, $E(\rho,v)=E_{J}(\rho)$ and $\rho$ locally minimizes $E_{J}(\rho)$ on ${R}\left(\mathbb{R}^{3}\right)$;
  3. (iii)
    $\rho$ is continuous on $\mathbb{R}^{3}$;
  4. (iv)
    on each connected component $\Omega_{i}$ of $\{\rho>0\},\rho$ satisfies
    $A^{\prime}(\rho(x))=\left[\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+V_{\rho}(x)+\lambda_{i}\right]_{+}$ (EL)
    for some Lagrange multiplier $\lambda_{i}<0$ depending on the component;
  5. (v)
    the equations (EL) continue to hold on a $\delta$-neighbourhood of the $\Omega_{i}$;
  6. (vi)
    if $P(\rho)$ satisfies assumption (F4), then $\rho\in C^{1}(\{\rho>0\})$;
  7. (vii)
    if $P(\rho)$ satisfies assumption (F4’), then $\rho$ satisfies the reduced Euler-Poisson equations (EP’), where the pressure $P(\rho)$ is a function of the density under assumptions (F1)(F2)(F3).
  8. (viii)
    this solution is stable with respect to $L^{\infty}$-small perturbations of the Lagrangian fluid variables.
Remark 2.26.
On each connected component $\Omega_{i}$ of $\{\rho>0\}$, let $W_{\rho}(x)=\frac{J^{2}}{2I^{2}(\rho)}r^{2}\left(x\right)+V_{\rho}\left(x\right)+\lambda_{i}$, then since $A^{\prime}$ is strictly increasing and $A^{\prime}(0)=0$, one can see $A^{\prime}(\rho(x))=W_{\rho}(x)$ on $\Omega_{i}$ due to (EL). (See also the proof of Proposition 5.8.)
Remark 2.27.
One might wonder whether $\rho$ in Theorem 2.25 (ii) is rotating. In fact, we can construct a truly rotating solution from it. Due to the discussion in subsection 2.2, if $\widetilde{\rho}$ is a local minimizer over ${R}\left(\mathbb{R}^{3}\right)$ of $E_{J}(\rho)$ — or by Corollary 4.9, even just over ${R}_{0}(\mathbb{R}^{3})$ — then we may define:
$({\rho}(t,x),v(t,x)):=\left(\widetilde{\rho}\left(R_{-wt}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$
where $\omega=\frac{J}{I\left(\rho\right)}$, then $({\rho}(t,x),v(t,x))$ gives a uniform rotating binary star system (or star-planet system) such that $E(\rho,v)$ and $E_{J}(\rho)$ is local minimum. Moreover, $({\rho}(t,x),v(t,x))$ solves (EP) with $V(t,x)=V_{\widetilde{\rho}}\left(R_{-\omega t}x\right)$.
Remark 2.28.
In McCann’s paper [31], he uses the terminology “chemical potential” to call the “Lagrange multiplier” $\lambda_{i}$ in Parts (iv) and (v) here. Although it seems no chemical reaction occurs when stars move or rotate or are at rest, the origin of the term “chemical potential” can trace back to “the Lagrange multiplier $\mu=\frac{\partial E}{\partial N}$ in the Euler–Lagrange variational equation for the ground state density as the chemical potential of the electrons” [7]. We notice by Part (ii) that $\rho$ is a minimizer on some set with some restriction. Therefore, another way to understand this term is to identify “Lagrange multiplier” with the derivative of energy by (generalized) Lagrange multiplier theorem Theorem A.4 and Remark A.5, thus it can be called “potential”. Just as the derivative of gravitational potential energy is referred to as gravitational potential (up to sign).
Remark 2.29.
Notice in Theorem 2.25, since the local minimizer $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$, we know $\bar{x}(\rho)=(0,0,0)^{T}$ in (EL). But when McCann tried to show the constrained minimizer $\rho^{J}$ (i.e. $J$ is fixed) is a Wasserstein $L^{\infty}$ local energy minimizer, he did not always require the center of mass of $\rho^{J}$ to be at original point. Therefore, we retain the term $\bar{x}(\rho)$ here to be consistent with the form of the Euler-Lagrange equations in McCann’s paper [31, Section 6].
Remark 2.30.
Thanks to $E(\rho,v)=E_{J}(\rho)$, we can further understand and elaborate on the meaning of Part (viii):
Let $\rho$ represents the original density profile, and $Y_{t}(y)\in\mathbb{R}^{3}$ represents the position of the fluid particles at time $t$ which originated at $Y_{0}(y)=y$, then the density profile $\rho_{t}$ after time $t$ satisfies ${Y_{t}}_{\#}\rho=\rho_{t}$, where $Y_{t}$$_{\#}$ is the push-forward operator given in Definition 4.1. From the definition of Wasserstein metric, we have
$W^{\infty}\left(\rho_{s},\rho_{t}\right)\leq\left\|Y_{s}-Y_{t}\right\|_{L^{\infty},\rho}$ (25)
Here $\|h\|_{L^{\infty}(S),v}$ denotes the supremum of $|h|$ over $S$, discarding sets of $v$-measure zero.
If we further assume the fluid particles move with bounded velocities, then $Y_{t}(y)$ will be a Lipshitz function of $t$ uniformly in $y$, and it is evident that (25) will be controlled by a multiple of |st||s-t|. Thus $\rho_{t}$ is a continuous function with respect to time $t$.
[…]
The last inequality comes from Proposition 2.17.
Those arguments suggest an $L^{\infty}$-small perturbation of the Lagrangian fluid variables produces only a $W^{\infty}$-small perturbation of the density and thereby a local energy minimum $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$ must be physically stable. We will introduce more related content in Section 4.
Remark 2.31.
Although Theorem 2.25 applies to energy minimizers subject only to a constraint on the z-component $J_{z}(\rho,v):=\hat{e}_{z}\cdot\boldsymbol{J}(\rho,v)$ of the angular momentum, it can be extended to the case of physical interest. To be precise, if $(\rho,v)$ minimizes $E(\rho,v)$ locally subject to the constraint $J_{z}(\rho,v)=J$, thanks to Theorem 2.25 (i)(ii), the angular momentum of $(\rho,v)$ is $\boldsymbol{J}(\rho,v)=J\hat{e}_{z}$, which means $(\rho,v)$ also minimizes $E(\rho,v)$ locally subject to the more restricted constraints $\boldsymbol{J}(\rho,v)=J\hat{e}_{z}$ [31, Corollary 2.2].
On the other hand, if the topology on ${V}\left(\mathbb{R}^{3}\right)$ is such that the map taking $\boldsymbol{w}\in\mathbb{R}^{3}$ to $v(x):=\boldsymbol{w}\times x\in{V}\left(\mathbb{R}^{3}\right)$ is continuous, then a converse to the statement above is also true: any local minimizer subject to the vector constraint $\boldsymbol{J}(\rho,v)=J\hat{e}_{z}$ also minimizes locally among the larger class of competitors with prescribed $J_{z}(\rho,v)=J$ [31, Remark 2.3]. This condition that $w\in\mathbb{R}^{3}\mapsto v(x):=w\times x\in{V}\left(\mathbb{R}^{3}\right)$ is continuous may initially appear to have no obvious use. But we can consider the following arguments: due to [31, Remark 3.3], we can have an similar result to 2.17: if $\boldsymbol{J}$ is prescribed, the velocity which minimizes the energy is given by $\boldsymbol{v}(\rho,\boldsymbol{J})(x)=\boldsymbol{w}(\rho,\boldsymbol{J})\times x$, where $\boldsymbol{w}(\rho,\boldsymbol{J})\in\mathbb{R}$ is the unique angular velocity compatible with the given density $\rho$ and angular momentum $\boldsymbol{J}$. Notice the axis $\boldsymbol{w}$ of rotation may not coincide with the $z$-axis. One can show when $\rho$ is sufficiently close to $\rho^{\prime}$ in the sense of Wasserstein $L^{\infty}$ distance, $w(\rho,\boldsymbol{J})$ is sufficiently close to $w^{\prime}(\rho^{\prime},\boldsymbol{J^{\prime}})$. This condition that $w\in\mathbb{R}^{3}\mapsto v(x):=w\times x\in{V}\left(\mathbb{R}^{3}\right)$ is continuous ensures that $\boldsymbol{v}(\rho,\boldsymbol{J})$ is sufficiently close to $\boldsymbol{v^{\prime}}(\rho^{\prime},\boldsymbol{J^{\prime}})$ in $V(\mathbb{R}^{3})$. In another word, local perturbations in ${R}_{0}\left(\mathbb{R}^{3}\right)$ can create local perturbations in ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$, which can help us to prove that any local minimizer subject to the vector constraint $\boldsymbol{J}(\rho,v)=J\hat{e}_{z}$ also minimizes locally among the larger class of competitors with prescribed $J_{z}(\rho,v)=J$, using similar arguments to [31, Remark 2.3 and Section 3].
Remark 2.32 (Equivalence between (EP’) and (EL)).
Now we can obtain a solution to (EP’) from the Euler-Lagrange equation (EL) thanks to Theorem 2.25 (iv). Conversely, for $\rho\in R_{0}(\mathbb{R}^{3})$ with N connected components to be a solution of (EP’), similar arguments as in subsection 2.2 shows that (EL) in Theorem 2.25 (iv) is necessary as we discuss in Remark 2.24.
By the arguments of Auchmuty and Beals [5] or Li [26] or McCann [31] we know a constrained energy minimizer on $W_{m,J}$ exists and its support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$. Then such constrained energy minimizer is indeed a local minimizer due to Lemma 4.7 (ii), and hence a solution to (EP’) after translation.
Theorem 2.33 (Existence of Binary Stars [31, Theorem 6.1, Corollary 6.2]).
[…]
What’s more, after another translation the center of mass of $\widetilde{\rho}$ can be 0 and is a local minimizer of $E_{J}(\rho)$. Let $v(x):=\omega\hat{e}_{z}\times x=\omega\left(-x_{2},x_{1},0\right)^{T}$, where $\omega=\frac{J}{I(\widetilde{\rho})}$, then the pair $(\widetilde{\rho},v)$ minimizes $E(\rho,v)$ locally on ${R}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$ (thus on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right))$ subject to the constraint $J_{z}(\rho,v)=J$ or $\boldsymbol{J}(\rho,v)=J\hat{e}_{z}$. $\widetilde{\rho}$ satisfies reduced Euler-Poisson equations (EP’). Moreover, the uniformly rotating fluid $({\rho}(t,x),v(t,x))$ solves (EP) with $V(t,x)=V_{\widetilde{\rho}}\left(R_{-\omega t}x\right)$, here $({\rho}(t,x),v(t,x))=\left(\widetilde{\rho}\left(R_{-\omega t}x\right),\omega\left(-x_{2},x_{1},0\right)^{T}\right)$.
3 Detailed Analysis of Theorem 2.25
We point out that in McCann’s proof of Theorem 2.25, the discussion on how to transition from (EL) to (EP’) is not particularly detailed. One can see intuitively from subsection 2.2 that a refined discussion on differentiability is required. Therefore, in this section, we will go through the proof of Theorem 2.25, give new ideas to establishing differentiability (Part (vi) and Part (vii) of Theorem 2.25), and explain the differences between our proof and McCann’s.
First we show the differentiability of $P\circ\phi=P\circ\left(A^{\prime}\right)^{-1}$, with the additional assumption (F4’).
Lemma 3.1 (Differentiability of $P\circ\phi$).
Assume $P$ satisfies (F1)(F2)(F3)(F4’), then $P\circ\phi$ is differentiable in $(0,\infty)$ and the derivative is $\phi(s)$. Moreover, $P\circ\phi$ is right differentiable at 0 with right derivatives $\phi(0)=0$. Here $\phi=\left(A^{\prime}\right)^{-1}$ exists by Remark 2.7.
Proof.
Given $s>0,h\in\mathbb{R}$ with $0<|h|\leq s$, let $t=\phi(s)$ and $t+l=\phi(s+h)$, then $h\rightarrow 0$ is equivalent to $l\rightarrow 0$ due to the continuity and monotonicity of $\phi$, and we also have $A^{\prime}(t)=s,A^{\prime}(t+l)=s+h$ since $\phi=\left(A^{\prime}\right)^{-1}$. Recall in Remark 2.5 we have $A^{\prime}(s)s-A(s)=P(s)$. Therefore, we have
$\frac{P\circ\phi(s+h)-P\circ\phi(s)}{h}$ $=\frac{P(t+l)-P(t)}{A^{\prime}(t+l)-A^{\prime}(t)}$
$=\frac{\left((t+l)A^{\prime}(t+l)-A(t+l)\right)-\left(tA^{\prime}(t)-A(t)\right)}{A^{\prime}(t+l)-A^{\prime}(t)}$
$=\frac{(t+l)\left(A^{\prime}(t+l)-A^{\prime}(t)\right)}{A^{\prime}(t+l)-A^{\prime}(t)}+\frac{lA^{\prime}(t)+A(t)-A(t+l)}{A^{\prime}(t+l)-A^{\prime}(t)}$
$=t+l+\frac{lA^{\prime}(t)+A(t)-A(t+l)}{A^{\prime}(t+l)-A^{\prime}(t)}$
Let $B(l)=lA^{\prime}(t)-A(t+l)$, then $B$ is differentiable and $B^{\prime}(l)=A^{\prime}(t)-A^{\prime}(t+l)\neq$ 0 if $l\neq 0$, since $A^{\prime}$ is strictly increasing. Notice $lA^{\prime}(t)+A(t)-A(t+l)=B(l)-B(0)$, by Cauchy’s mean value theorem [35, 5.9 Theorem], $\exists\theta\in(0,1)$, such that $(B(l)-B(0))A^{\prime\prime}(t+\theta l)=\left(A^{\prime}(t+l)-A^{\prime}(t)\right)B^{\prime}(\theta l)$. Since $l\neq 0$, $\left(A^{\prime}(t+l)-A^{\prime}(t)\right)\neq 0$, $B^{\prime}(\theta l)\neq 0$, then both (B(l)B(0))(B(l)-B(0)) and $A^{\prime\prime}(t+\theta l)$ are also not 0. Then
$\frac{lA^{\prime}(t)+A(t)-A(t+l)}{A^{\prime}(t+l)-A^{\prime}(t)}$ $=\frac{(B(l)-B(0))B^{\prime}(\theta l)}{(B(l)-B(0))A^{\prime\prime}(t+\theta l)}$
$=\frac{A^{\prime}(t)-A^{\prime}(t+\theta l)}{A^{\prime\prime}(t+\theta l)}$
By Lagrange’s mean value theorem [35, 5.8 Theorem], $\exists\widetilde{\theta}\in(0,1)$, such that $A^{\prime}(t)-A^{\prime}(t+\theta l)=(-\theta l)A^{\prime\prime}(t+\widetilde{\theta}\theta l)$. Notice $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}$ when $s>0$. Thus
$\frac{A^{\prime}(t)-A^{\prime}(t+\theta l)}{A^{\prime\prime}(t+\theta l)}$ $=\frac{(-\theta l)A^{\prime\prime}(t+\widetilde{\theta}\theta l)}{A^{\prime\prime}(t+\theta l)}$
$=-\frac{\theta l(t+\theta l)}{t+\widetilde{\theta}\theta l}\cdot\frac{P^{\prime}(t+\widetilde{\theta}\theta l)}{P^{\prime}(t+\theta l)}$
By Taylor’s theorem [35, 5.15 Theorem], we know $\exists\theta_{1}\in(0,1),\theta_{2}\in(0,1)$, $P^{\prime}(t+\theta l)=P^{\prime}(t)+\theta lP^{\prime\prime}(t)+\frac{(\theta l)^{2}}{2}P^{(3)}(t)+\cdots+\frac{(\theta l)^{n}}{n!}P^{(n+1)}\left(t+\theta_{1}\theta l\right)=\frac{(\theta l)^{n}}{n!}P^{(n+1)}$ $\left(t+\theta_{1}\theta l\right)\neq 0$ due to (F4’). Similarly, we have $P^{\prime}(t+\widetilde{\theta}\theta l)=\frac{(\widetilde{\theta}\theta l)^{n}}{n!}P^{(n+1)}\left(t+\theta_{2}\widetilde{\theta}\theta l\right)\neq 0$. Thus $\left|\frac{P^{\prime}(t+\widetilde{\theta}\theta l)}{P^{\prime}(t+\theta l)}\right|\leq 2$ when $|l|$ small enough. As $l\rightarrow 0$, we have $\frac{A^{\prime}(t)-A^{\prime}(t+\theta l)}{A^{\prime\prime}(t+\theta l)}\rightarrow 0$, and then $\frac{P\circ\phi(s+h)-P\circ\phi(s)}{h}\rightarrow t=\phi(s)$ as $h\rightarrow 0$. Thus, the derivative of $P\circ\phi$ is $\phi(s)$. By the same arguments we know $P\circ\phi$ is right differentiable at 0 with right derivatives $\phi(0)=$ 0. ∎
Remark 3.2.
Since (F4) implies (F4’), thus the result in Lemma 3.1 also holds true if $P$ satisfies (F1)(F2)(F3)(F4). One can also apply Cauchy’s mean value theorem directly to $\frac{P\circ\phi(s+h)-P\circ\phi(s)}{h}=\frac{P(t+l)-P(t)}{A^{\prime}(t+l)-A^{\prime}(t)}$ and obtain $\frac{P\circ\phi(s+h)-P\circ\phi(s)}{h}=\frac{P(t+l)-P(t)}{A^{\prime}(t+l)-A^{\prime}(t)}=\frac{P^{\prime}(t+\theta l)}{A^{\prime\prime}(t+\theta l)}=t+\theta l\rightarrow t=\phi(s)$ as $h\rightarrow 0$. We can do this because (F4), together with $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}$ if $s>0$, implies $A^{\prime\prime}(t+\theta l)\neq 0$.
We skip the proof of Part (i) and Part (viii) of Theorem 2.25, one can refer to McCann’s paper [31]. Part (ii) and Part (iii) are closely related to the context, so we revisit McCann’s proof here. The other parts of the proof are relatively new, either addressing details not mentioned by McCann (Part (iv-v)) or providing new approachs to proving differentiability and explaining the necessity of modifying the original statements of the conclusions (Part (vi-vii)).
Proof of Part (ii-vi) of Theorem 2.25.
Part (ii) essentially follows from Corollary 2.18 and Corollary 4.9. We note that $E(\rho,v)=E_{J}(\rho)$ due to the definition of $v(x):=\frac{J\hat{e}_{z}\times x}{I(\rho)}$, $E$ (6), $E_{J}$ (16), $T$ (9), and $T_{J}$ (22).
Due to Part (ii), it suffices to consider $\rho$ that locally minimizes $E_{J}(\rho)$ on $R(\mathbb{R}^{3})$. We will give proofs of “weak” versions of Parts (iv-v) (Proposition 5.8, Proposition 5.11 and Remark 5.12) later in subsection 5.2, that is, (EL) holds for almost all points in the region we consider with negative $\lambda_{i}$. Based on this “weak” version of (v), we obtain the global continuity in Part (iii) as the following: first notice that almost everywhere in the region $\{\rho>0\}$, (EL) holds true and $A^{\prime}(\rho)>0$ implies $\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x)+V_{\rho}(x)+\lambda_{i}>0$, thus on $\{\rho>0\}$, $A^{\prime}(\rho(x))=\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x)+V_{\rho}(x)+\lambda_{i}\leq V_{\rho}(x)+C$ since $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$ has compact support. This result, together with Proposition A.7 (Hardy-Littlewood-Sobolev Inequality) or Proposition A.9, can help to show the bound of $\|\rho\|_{L^{p}}$ and $\left\|V_{\rho}\right\|_{L^{\widetilde{p}}}$ with $p$ and $\widetilde{p}$ increasing alternately and reaching infinity in finite steps. (Bootstrap method, similar as [5, Lemma 3 and Theorem A]). Then $V_{\rho}$ is continuously differentiable thanks to Proposition A.11. Then on a $\delta$-neighbourhood of $\{\rho>0\}$, denoted by $\Omega_{\delta}$, which contains $\partial\{\rho>0\}$, we have $\rho=\phi\circ\left[W_{\rho}\right]_{+}$ a.e., where $W_{\rho}(x)=\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x)+V_{\rho}(x)+\lambda_{i},\phi=\left(A^{\prime}\right)^{-1}$ exists and continuous by Remark 2.7. It means $\rho$ has a continuous representative there. Notice the points in $\overline{\mathbb{R}^{3}\backslash\Omega_{\delta}}$ are the interior points of $\{\rho=0\}$, hence $\rho$ is also continuous in $\overline{\mathbb{R}^{3}\backslash\Omega_{\delta}}$. Thus, we conclude Part (iii).
Due to the global continuity of $\rho$ and $V_{\rho}$ (that is, they are continuous on $\mathbb{R}^{3}$), we can come back to strengthen the “weak” versions of Parts (iv-v), replace “almost all points” by “all points” and get Parts (iv) and Part (v).
For Part (vi), notice in the region $\{\rho>0\}$, we have $\rho=\phi\circ W_{\rho}$ by Remark 2.26. Thanks to Part (iii) and $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$, we know $\rho\in C_{c}^{0}\left(\mathbb{R}^{3}\right)$, where $C_{c}^{0}\left(\mathbb{R}^{3}\right)$ denotes the family of continuous functions with compact support. In particular, $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{q}\left(\mathbb{R}^{3}\right)$ for any $q>3$. By Proposition A.11, we know $V_{\rho}$ is continuous differentiable in $\mathbb{R}^{3}$. Since $P(\rho)$ satisfies assumption (F4), by Remark 2.8 we know $\phi\in C^{1}((0,\infty))$. Apply chain rule (one can refer to [38, Theorem 10.1.15]), we obtain $\rho\in C^{1}(\{\rho>0\})$. ∎
Remark 3.3.
Theorem 2.25 (vi) is slightly different from the one in McCann’s theorem [31, Theorem 2.1 (vi)], which states “where $\rho$ is positive, it has as many as derivatives as the inverse of $A^{\prime}(\rho)$”. In McCann’s proof, he said “Because $V_{\rho}$ gains a derivative from $\rho$, smoothness of $\rho$ where positive follows from a bootstrap in (15)”, where “(15)” is (EL) in this paper. It is true that once $\rho\in W^{k,q}\left(\mathbb{R}^{3}\right)$ with some $q>3$ then we can obtain $V_{\rho}\in W^{k+1,\infty}\left(\mathbb{R}^{3}\right)\cap C^{k+1}\left(\mathbb{R}^{3}\right)$ by the similar arguments in Proposition A.11. But so far we only know $\rho\in C^{1}\left(\left\{{\rho>0}\right\}\right)\cap C_{c}^{0}\left(\mathbb{R}^{3}\right)$, one thing that is not clear is whether $\rho\in W^{1,q}\left(\{\rho>0\}\right)$, and hence whether $\rho\in W^{1,q}\left(\mathbb{R}^{3}\right)$. The difficulty lies in the boundary behaviour: although $\rho$ is globally continuous and vanishes on $\partial\{\rho>0\}$, we do not know the regularity of $\partial\{\rho>0\}$ and we cannot rule out the possibility that $\nabla\rho$ blows up near the boundary. It means we might not be able to apply the result of trace operator (e.g., [25, Exercise 15.26]) to glue $\rho\cdot\mathbf{1}_{\{\rho>0\}}$ and $\rho\cdot\mathbf{1}_{\{\rho=0\}}$ (here $\mathbf{1}_{E}$ denotes the indicator function of $E$).
If we do not know whether $\rho\in W^{1,q}\left(\mathbb{R}^{3}\right)$, we might also not know whether $V_{\rho}\in W^{2,\infty}\left(\mathbb{R}^{3}\right)\cap C^{2}\left(\mathbb{R}^{3}\right)$. Even if we assume, for example, $\phi\in C^{2}\left((0,\infty)\right)$, we know $\rho(x)=\phi\circ W_{\rho}(x)=\phi\circ\left(\frac{J^{2}}{2I^{2}(\rho)}r^{2}\left(x\right)+V_{\rho}(x)+\lambda_{i}\right)$ on $\Omega_{i}$ due to Remark 2.26, but it might not be enough to show $\rho\in C^{2}(\{\rho>0\})$ if we do not know $V_{\rho}\in W^{2,\infty}\left(\mathbb{R}^{3}\right)\cap C^{2}\left(\mathbb{R}^{3}\right)$. Therefore, it is not quite convincing if $\phi\in C^{r}((0,\infty))$ implies $\rho\in C^{r}(\{\rho>0\})$ if $r\geq 2$.
Remark 3.4.
Theorem 2.25 (vi) suggests $\phi=\left(A^{\prime}\right)^{-1}\in C^{1}((0,\infty))$ implies $\rho\in C^{1}(\{\rho>0\})$. But whether the converse is correct needs more discussion. For example, assume $A^{\prime\prime}(s)=0$ for some $s>0$, then the “derivative” of $\phi$ at point $t:=A^{\prime}(s)$ should be $\infty$, or more precisely, $\lim\limits_{h\rightarrow 0}\frac{\phi(t+h)-\phi(t)}{h}=\infty$. To show it, one can following the proof of inverse function theorem [38, Theorem 10.4.2] and notice $A^{\prime}$ and $\phi$ are strictly increasing.
Assume there is a point $x_{0}\in\{\rho>0\}$ such that $W_{\rho}(x_{0})=\frac{J^{2}}{2I^{2}(\rho)}r^{2}\left(x_{0}\right)+V_{\rho}\left(x_{0}\right)+\lambda_{i}=t$, i.e. $\phi\circ W_{\rho}(x_{0})=\rho(x_{0})$ due to Remark 2.26, one may guess the gradient of $\rho$ does not exist at $x_{0}$, but could have problem if $\nabla W_{\rho}=\nabla\left(\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x)+V_{\rho}(x)+\lambda_{i}\right)$ happens to be 0 at $x_{0}$. A trivial example which helps to describe such situation is $h(s)=g\circ f(s)$, where $g(t)=\left\{\begin{array}[]{ll}-\sqrt{-t},&t<0\\ \sqrt{t},&t\geq 0\end{array},f(s)=0\right$.. Then $h(s)=0$ has derivative at $s_{0}=0$, although the “derivative” of $g$ at $t_{0}=f\left(s_{0}\right)=$ 0 is $\infty$. Therefore, it is not obvious whether $\rho\in C^{1}(\{\rho>0\})$ can implies $\phi=\left(A^{\prime}\right)^{-1}\in C^{1}((0,\infty))$.
Fortunately, when we try to show $\rho$ satisfies (EP’), we do not necessarily need to deal with the gradient of $\rho$; we only need to consider the gradient of $P(\rho)$. As a result, (F4) can be relaxed to (F4’). Under the conditions of (F4’), we can still prove gradient of $P(\rho)$ exists and then $\rho$ satisfies (EP’), as shown in the proof of Part (vii) of Theorem 2.25 below.
Remark 3.5.
It’s important to note that it’s because $\rho$ is a $W^{\infty}$-local energy minimizer that we can extend (EL) to a neighborhood of $\{\rho>0\}$, and then obtain the global continuity Part (iii). It helps to exclude the case that minimizer (star) has non-zero pressure on the boundary. Moreover, this will be crucial in showing the existence of $\nabla P(\rho)$ on the boundary $\partial\{\rho>0\}$, and then Part (vii) in the following.
Proof of Part (vii) of Theorem 2.25.
$\rho\in C_{c}^{0}\left(\mathbb{R}^{3}\right)$ implies $\nabla V_{\rho}$ exists and is continuous in $\mathbb{R}^{3}$ thanks to Proposition A.11, so is $\nabla W_{\rho}$, where $W_{\rho}(x)=\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x)+V_{\rho}(x)+\lambda_{i}$ as in the proof of Part (iii-vi). In the region $\{\rho>0\}$, (EL) holds true and $A^{\prime}(\rho)>0$ implies $W_{\rho}>0$. $\nabla W_{\rho}$ exists means $\nabla A^{\prime}(\rho)$ exists. Thanks to Lemma 3.1, we know
$\nabla(P\circ\rho)(x)$ $=\nabla\left(P\circ\phi\circ W_{\rho}\right)(x)$
$=(P\circ\phi)^{\prime}\left(W_{\rho}(x)\right)\nabla W_{\rho}(x)$
$=\phi\left(W_{\rho}(x)\right)\nabla W_{\rho}(x)$
$=\phi\left(W_{\rho}(x)\right)\nabla A^{\prime}(\rho(x))$
$=\rho(x)\nabla A^{\prime}(\rho(x))$
Then we take the gradient of (EL) in $\{\rho>0\}$. By the similar computations in subsection 2.2, we have (EP’) holds true in $\{\rho>0\}$.
In the interior of $\{\rho=0\}$, easy to check (EP’) still holds true since $\nabla P(\rho(x))$ exists and $\nabla P(\rho(x))=0$, and $\nabla V_{\rho}(x)$ exists and finite.
The points remain to be checked is those on the boundary $\partial\{\rho=0\}=\partial\{\rho>0\}$. Since $\rho$ is continuous, we know $\rho=0$ on $\partial\{\rho=0\}$ and then similarly as above we have $\omega^{2}\rho(x)P_{12}(x)=-\rho(x)\nabla V_{\rho}(x)=0$. But to show (EP’) holds, we still need to check if $\nabla P(\rho(x))$ exists and is 0 on $\partial\{\rho=0\}$.
Given $x_{0}\in\partial\{\rho=0\}$ and $\left\{x_{n}\right\}\subset\mathbb{R}^{3}$, where $x_{n}=x_{0}+h_{n}\widehat{e_{1}},\widehat{e_{1}}=(1,0,0)^{T},h_{n}\neq 0$, and $\lim\limits_{n\rightarrow\infty}h_{n}=0$. If $\rho\left(x_{n}\right)=0$ for all $n$, then $P\left(\rho\left(x_{n}\right)\right)=P\left(\rho\left(x_{0}\right)\right)=0$, and $\lim\limits_{n\rightarrow\infty}\frac{P\left(\rho\left(x_{n}\right)\right)-P\left(\rho\left(x_{0}\right)\right)}{h_{n}}=0$. If $\rho\left(x_{n}\right)>0$ for all $n$, then $y_{n}:=W_{\rho}\left(x_{n}\right)>0$ and $\rho\left(x_{n}\right)=\phi\left(y_{n}\right)$ as above.
We first claim $x_{0}\in\partial\{\rho>0\}$ implies $y_{0}:=W_{\rho}\left(x_{0}\right)=0$ by continuity. In fact, $y_{0}>0$ would imply $\rho\left(x_{0}\right)>0$, while $y_{0}<0$ would imply $x_{0}$ is in the interior of $\{\rho=0\}$, and both cases cannot happen since $x_{0}\in\partial\{\rho>0\}$. In particular, $\rho(x_{0})=\phi(y_{0})=0$.
Then, by Lemma 3.1 and Lagrange’s mean value theorem [35, 5.8 Theorem], $\exists\theta_{1}\in(0,1),\theta_{2}\in(0,1)$, such that
$\frac{P\left(\rho\left(x_{n}\right)\right)-P\left(\rho\left(x_{0}\right)\right)}{h_{n}}$ $=\frac{P\left(\phi\left(y_{n}\right)\right)-P\left(\phi\left(y_{0}\right)\right)}{y_{n}-y_{0}}\cdot\frac{y_{n}-y_{0}}{h_{n}}$
$=\phi\left(y_{0}+\theta_{2}\left(y_{n}-y_{0}\right)\right)\cdot\partial_{x_{1}}W_{\rho}\left(x_{0}+\theta_{1}\left(x_{n}-x_{0}\right)\right)$
$\lim\limits_{n\rightarrow\infty}x_{n}=x_{0}$ implies $\left\{x_{n}\right\}$ is bounded, so is $\left\{y_{n}\right\}$. Therefore, $\{\partial_{x_{1}}W_{\rho}(x_{0}+\theta_{1}(x_{n}-x_{0}))\}$ is a bounded sequence. $\lim\limits_{n\rightarrow\infty}\phi\left(y_{0}+\theta_{2}\left(y_{n}-y_{0}\right)\right)=\phi\left(y_{0}\right)=0$ implies $\lim\limits_{n\rightarrow\infty}\frac{P\left(\rho\left(x_{n}\right)\right)-P\left(\rho\left(x_{0}\right)\right)}{h_{n}}=0$.
Now we have considered the cases $\rho\left(x_{n}\right)=0$ and $\rho\left(x_{n}\right)>0$ and show $\lim\limits_{n\rightarrow\infty}\frac{P\left(\rho\left(x_{n}\right)\right)-P\left(\rho\left(x_{0}\right)\right)}{h_{n}}=0$ in both cases. It is a sufficient condition to show $\partial_{x_{1}}P(\rho)$ exists and is 0 at $x_{0}$. Otherwise, we could have a sequence $\left\{\widetilde{h_{n}}\right\}$ converging to 0, $\widetilde{x_{n}}=x_{0}+\widetilde{h_{n}}\widehat{e_{1}},\left\{\widetilde{x_{n}}\right\}$ (up to subsequence) satisfies either $\rho\left(\widetilde{x_{n}}\right)>0$ for all $n$ or $\rho\left(\widetilde{x_{n}}\right)=0$ for all $n$, $\frac{P\left(\rho\left(\widetilde{x_{n}}\right)\right)-P\left(\rho\left(\widetilde{x_{0}}\right)\right)}{h_{n}}$ would not converge to 0. But the result above shows $\lim\limits_{n\rightarrow\infty}\frac{P\left(\rho\left(\widetilde{x_{n}}\right)\right)-P\left(\rho\left(\widetilde{x_{0}}\right)\right)}{h_{n}}=0$, which leads to a contradiction. Similarly, $\partial_{x_{2}}P(\rho)$ and $\partial_{x_{3}}P(\rho)$ also exist and are 0 at $x_{0}$. Therefore, $\nabla P(\rho(x))=0$ on $\partial\{\rho>0\}$ and then (EP’) holds true on $\partial\{\rho>0\}$, which gives us Part (vii). ∎
Remark 3.6.
Theorem 2.25 (vii) is also slightly different from the one in McCann’s theorem [31, Theorem 2.1 (vii)], which states “If $P(\rho)$ is continuously differentiable on $0,\infty)$, then $\rho$ satisfies (EP’)”. Although there is only two-line proof of Part (vii) in McCann’s paper [31, Theorem 2.1]: “If $P(\rho)$ is continuously differentiable, then $A^{\prime\prime}(\rho)=\frac{P^{\prime}(\rho)}{\rho}$ and (vii) follows by taking the gradient of (15)”, where “(15)” is (EL) in this thesis, it is not quite trivial since we do not know whether $\nabla P(\rho)$ truly exists in $\mathbb{R}^{3}$ ((EL) holds only in a $\delta$-neighborhood). Moreover, even if $P$ is differentiable, the identity $A^{\prime\prime}(s)=\frac{P^{\prime}(s)}{s}$ for a scalar $s$ does not immediately imply $\nabla A^{\prime}(\rho)=\frac{\nabla P(\rho)}{\rho}$ for a function $\rho$. Indeed, if $\nabla\rho$ does not exist, expressions such as $\nabla A^{\prime}(\rho)=A^{\prime\prime}(\rho)\nabla\rho$ or $\frac{\nabla P(\rho)}{\rho}=\frac{P^{\prime}(\rho)\nabla\rho}{\rho}$ are not justified, let alone the possibility of canceling $\nabla\rho$ on both sides. We need more detailed discussions as above, under the assumption (F4’).
Remark 3.7.
When we show $\nabla P(\rho(x))$ exists on $\partial\{\rho>0\}$, one might want to follow the same arguments (chain rule) as in $\{\rho>0\}$, but there could be a problem that we might not have $\rho=\phi\circ W_{\rho}$ in a $\epsilon$-neighborhood of $\partial\{\rho>0\}$. However, it can still inspire us to consider the case $\rho\left(x_{n}\right)>0$ as above and follow the proof of chain rule.
Remark 3.8.
If (F4) holds, as we discuss in Remark 2.8, due to Part (iv) we know in the region $\{\rho>0\}$, $\nabla\rho$ exists, then $\nabla P(\rho)$ also exists and satisfies $\nabla P(\rho)=P^{\prime}(\rho)\nabla\rho=\rho A^{\prime\prime}(\rho)\nabla\rho=\rho\nabla A^{\prime}(\rho)$. This is what we show above, but the arguments here are easier, with this stronger assumption (F4), though such arguments may not be valid when we discuss the region $\{\rho=0\}$ or $\partial\{\rho>0\}$.
Remark 3.9.
Going through the proof, we can see Theorem 2.25 also holds true when the number of components $n$ is not 2. However, it may not be clear so far if such $W^{\infty}$ local minimizers exist or not when $n\geq 3$. Jang and Seok [23, 1. Introduction] say “when $N\geq 3$, uniformly rotating N-body stellar objects do not retain a variational characterization analogous to the binary case and they are not expected to be stable in general”.
Remark 3.10.
In Remark 2.14 and Theorem 2.33, we mention that the support of constrained minimizer lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$ implies the constrained minimizer is local minimizer hence satisfies Theorem 2.25. In fact, if we revisits the proof of Theorem 2.25, those results (excluding the equivalence result that $(\rho,v)$ minimizes $E(\rho,v)$ locally if and only if $\rho$ minimizes $E_{J}(\rho)$ locally) are also true without specifying it is a local minimizer. The crucial thing is deriving (EL) in a neighborhood of the support of constrained minimizer, which is true if the support lies in the interior of $\Omega_{m}\cup\Omega_{1-m}$. Then we can verify (EP’) at the boundary of the support.
4 Wasserstein $L^{\infty}$ Metric on ${R}\left(\mathbb{R}^{3}\right)$
In this section we revisit some basic concepts and properties of Wasserstein $L^{\infty}$ metric ($W^{\infty}$ metric), which is important to show the existence of solutions to (EP). Additionally, we present a new result stating that, under the topology induced by $W^{\infty}$ metric, given $\rho$, one can always find a $\sigma$ in the neighborhood of $\rho$ such that $\sigma$ lies in $L^{\infty}$. This result is helpful in Section 5, where we establish the existence of variational derivatives for local minimizers.
Let us first discuss the motivation of introducing $W^{\infty}$ metric: as McCann says in [31], when we discuss what local minima are, we should first specify the topology. The choice will be delicate. In order to have local minima, we hope the topology to be strong enough to preclude tunneling of mass (counterexample can be seen in [31, Example 3.6 and Remark 3.7]). On the other hand, to make the local minima physically meaningful, we hope the topology to be weak enough so that the evolution of physical flows is continuous (see Remark 2.30 or the text below). Thanks to some probability literatures, we can find a topology with these properties, that is, Wasserstein $L^{\infty}$ metric described e.g. by McCann [31], Ambrosio, Brué and Semola [2] or Givens and Shortt [18].
In a metric space $(X,d),\mathcal{B}(X)$ denotes its Borel $\sigma$-algebra and $\mathcal{M}(X)$ the set of the $\sigma$-additive functions $\tilde{\mu}:\mathcal{B}(X)\rightarrow\mathbb{R}$. Furthermore, we define the nonnegative measure space as:
$\mathcal{M}_{+}(X):=\{\tilde{\mu}\in\mathcal{M}(X):\tilde{\mu}\geq 0\}$
And define the probability (measure) space as:
$\mathcal{P}(X)=\left\{\tilde{\mu}\in\mathcal{M}_{+}(X):\tilde{\mu}(X)=1\right\}$
Definition 4.1 (Push Forward Measure).
Given a Borel function $f:X\rightarrow Y$, we define the push forward operator $f_{\#}:M(X)\rightarrow M(Y)$ by
$f_{\#}\tilde{\mu}(B):=\tilde{\mu}\left(f^{-1}(B)\right)\quad\forall B\in\mathcal{B}(Y)$
And call $f_{\#}\tilde{\mu}$ push forward measure.
Let $(X,d)$ be a metric space and $(S,\Sigma,v)$ be a probability space, given $\rho,\kappa\in\mathcal{P}(X)$, the Wasserstein $L^{\infty}$ distance between $\rho$ and $\kappa$ is defined as
$W^{\infty}(\rho,\kappa):=\inf\left\{\begin{array}[]{l}\|d(f(x),g(x))\|_{L^{\infty},v}\mid f:S\rightarrow X\text{ Borel },\\ f_{\#}v=\rho\text{ and }g:S\rightarrow X\text{ Borel, }g_{\#}v=\kappa\end{array}\right\}$ (26)
Here $\|h\|_{L^{\infty},v}$ denotes the supremum of $|h|$ over $S$, discarding sets of $v$-measure zero. We call $f,g$ the transport maps. Thanks to Strassen’s Theorem, one can check $W^{\infty}$ is truly a metric, as explained in Givens and Shortt [18].
Remark 4.2.
One might wonder whether we could find such transport maps. It turns out at least in some cases we can always find them. See for example [2, Theorem 1.11, Theorem 5.2].
Remark 4.3.
It turns out $W^{\infty}(\rho,\kappa)$ is actually independent of the probability space $(S,\Sigma,v)$ we choose. Due to this fact, we can choose $S=[0,1]$ with Lebesgue measure, or $(S,\Sigma,v)=(X,\mathcal{B}(X),\rho)$. If we consider the latter case, we have
$W^{\infty}(\rho,\kappa)$ $=\inf\left\{\begin{array}[]{l}\|d(f(x),g(x))\|_{L^{\infty},\rho}\mid f:X\rightarrow X\text{ Borel},\\ f_{\#}\rho=\rho\text{ and }g:X\rightarrow X\text{ Borel, }g_{\#}\rho=\kappa\end{array}\right\}$ (27)
$\leq\inf\left\{\|d(x,g(x))\|_{L^{\infty},\rho}\mid g:X\rightarrow X\text{ Borel},g_{\#}\rho=\kappa\right\}$
The right side of the inequality can be viewed as a (generalized) Monge’s formulation of the optimal transport problem (see e.g. [2, Section 1.2 and Section 8.1]).
Remark 4.4.
Notice in [2, Section 8.1], the authors define another Wasserstein $L^{\infty}$ distance (with an abuse of notation denoted by $\widetilde{W}^{\infty}$) via transport plans and Kantorovich’s formulation of the optimal transport problem. In general, we also have $\widetilde{W}^{\infty}(\rho,\kappa)\leq\inf\left\{\|d(x,g(x))\|_{L^{\infty}(\rho)}\mid g:X\rightarrow X\text{ Borel},g_{\#}\rho=\kappa\right\}$ [2, Section 2.2]. Moreover, when we consider Wasserstein $L^{p}$ distance given in [2], $p<\infty$, under additional assumption(s) if needed, we have $\widetilde{W}^{p}(\rho,\kappa)=(\inf\{\int_{X}d^{p}(x,g(x))d\rho(x)\mid g:X\rightarrow X\text{ Borel},g_{\#}\rho=\kappa\})^{\frac{1}{p}}$ [2, Section 2.2 or Section 5.1]. However, the link between $\widetilde{W}^{\infty}$ and $W^{\infty}$ seems not clear. In the following, our arguments will be based on $W^{\infty}$’s definition.
Remark 4.5.
$\rho\in\mathcal{P}(X)$ may be represented in many ways such as the probability distribution. Consider $X=\mathbb{R}^{3}$ with $d(x,y)=|x-y|,\forall x,y\in\mathbb{R}^{3},\rho$ can be viewed as a law of a vector-valued random variable $f:S\rightarrow\mathbb{R}^{3}$ on the probability space $(S,\Sigma,v)$, with $f_{\#}v=\rho$. And, $W^{\infty}(\rho,\kappa)$ may now be defined as an infimum over all random variable representations of $\rho$ and $\kappa$ on a space $(S,\Sigma,v)$.
Remark 4.6.
Since the functions in ${R}\left(\mathbb{R}^{3}\right)$ have unit mass, we can view them as measures. That is, given $\rho\in{R}\left(\mathbb{R}^{3}\right)$, define $\widetilde{\rho}$ as the following: $\forall B\in\mathcal{B}\left(\mathbb{R}^{3}\right),\widetilde{\rho}(B):=\int_{B}\rho$. One can check in this definition, $\widetilde{\rho}\in\mathcal{P}\left(\mathbb{R}^{3}\right),\widetilde{\rho}$ is absolutely continuous with respect to $\mu$ (denoted by $\widetilde{\rho}\ll\mu$), where $\mu$ is Lebesgue measure, and the derivative (density) of $\widetilde{\rho}$ with respect to $\mu$ is $\rho$ [16, Section 1.6]. With an abuse of notation, we denote $\widetilde{\rho}$ by $\rho$ as well. Therefore, we can define the Wasserstein $L^{\infty}$ distance between two functions in ${R}\left(\mathbb{R}^{3}\right)$. Note that although $W^{\infty}(\rho,\kappa)$ may be infinite on ${R}\left(\mathbb{R}^{3}\right)$, it is finite whenever $\rho$ and $\kappa$ are of bounded support by Lemma 4.7 (i) below.
As McCann explained [31, Section 5], it turns out the Wasserstein $L^{\infty}$ metric is not unphysically strong: for instance, in the Lagrangian description of fluid mechanics [6, Section 1.2] [39] [24, Chapter 1], the state of a fluid system is specified by its original density profile $\rho\in{R}\left(\mathbb{R}^{3}\right)$, together with the positions of the fluid particles as a function of time. In Remark 2.30, we already see $\rho_{t}\in{R}\left(\mathbb{R}^{3}\right)$, as a function of time, evolves continuously with respect to the topology induced by Wasserstein $L^{\infty}$ metric, at least for bounded fluid velocities. Here ${Y_{t}}_{\#}\rho=\rho_{t}$, and $Y_{t}(y)\in{R}\left(\mathbb{R}^{3}\right)$ represents the position of the fluid at time $t$ which originated at $Y_{0}(y)=y$. Therefore, a local energy minimum $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$ is physically stable.
We notice that the definition of Wasserstein distance can be generated to two functions in $t{R}\left(\mathbb{R}^{3}\right)=\left\{\left.\rho\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\right\rvert\,\rho\geq 0,\int\rho=t\right\}$, with $t\neq 0$. Moreover, one can observe that given $\rho,\sigma$ in ${R}\left(\mathbb{R}^{3}\right),W^{\infty}(t\rho,t\sigma)=W^{\infty}(\rho,\sigma)$ (or $\widetilde{W}^{\infty}(t\rho,t\sigma)=\widetilde{W}^{\infty}(\rho,\sigma))$, which is different from Wasserstein $L^{p}$ distance.
We give some elementary properties required of $W^{\infty}$. The first 5 properties essentially are picked from [31, Lemma 5.1]. Here $\operatorname{spt}(\rho-\kappa)\subset\mathbb{R}^{3}$ denotes the support of the signed measure $\rho-\kappa$, while a $\delta$-neighbourhood is defined as in 2.1 (viii).
Lemma 4.7 (Simple Properties of the Wasserstein $L^{\infty}$ Metric).
Let $\rho,\kappa$ in $t{R}\left(\mathbb{R}^{3}\right)$, then
  1. (ii)
    $W^{\infty}(\rho,\kappa)$ does not exceed the diameter of $\operatorname{spt}(\rho-\kappa)$;
  2. (ii)
    if $W^{\infty}(\rho,\kappa)<\delta$, each connected component of the $\delta$-neighbourhood of spt $\rho$ has the same mass for $\kappa$ as for $\rho$;
  3. (iii)
    $W^{\infty}\left(\rho,f_{\#}\rho\right)\leq\|f-id\|_{\infty,\rho}$ for $f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ measurable and $id(x):=x$;
  4. (iv)
    the centers of mass satisfy $|\bar{x}(\rho)-\bar{x}(\kappa)|\leq W^{\infty}(\rho,\kappa)$;
  5. (v)
    the moment of inertia $I(\rho)$ depends continuously on $\rho$;
  6. (vi)
    given $\rho\in t{R}\left(\mathbb{R}^{3}\right),\forall\epsilon>0,\exists\sigma\in t{R}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$, such that $W^{\infty}(\rho,\sigma)<\epsilon$. Moreover, if spt $\rho$ is bounded, then spt $\sigma$ is also bounded.
Proof.
(ii-$v$) comes from [31, Lemma 5.1] with replacing ${R}\left(\mathbb{R}^{3}\right)$ by $t{R}\left(\mathbb{R}^{3}\right)$, which proofs are from the definition of $W^{\infty}$; the idea behind part (iii) can also be seen in Remark 2.30.
To prove ($vi$), the intuitive strategy is to “cut off” the function values of $\rho$ at the region where the value is relatively large, and then redistribute the value cut off to surrounding area where the values are relatively smaller, and then prove that the $W^{\infty}$ distance between the rearranged function and the original function is very small. To be precise, we carry out the following procedure. Without loss of generality we assume $t=1$, and we first divide the space $\mathbb{R}^{3}$ as $\mathbb{R}^{3}=\left(\cup_{n=1}^{\infty}C\left(x_{n}\right)\right)\cup D$. Here $\left\{C\left(x_{n}\right)\right\}$ is a family of (countably) many disjoint small open cubes, each with a diagonal length equal to $\frac{\epsilon}{2}$, volume equal to $V:=\left(\frac{\epsilon}{2\sqrt{3}}\right)^{3}$, centered at $x_{n}$. $D$ is the union of the remaining boundaries, $0\in D$, and $\mu(D)=0$. Since $\int\rho=1=\sum_{N=0}^{\infty}\int_{\{x\mid N\leq\rho(x)<N+1\}}\rho,\exists R>1$, such that $\int_{\{x\mid\rho(x)>R\}}\rho<\frac{V}{4}$. Let $\rho_{n}=\rho\cdot\mathbf{1}_{C\left(x_{n}\right)}$, where $\mathbf{1}_{C\left(x_{n}\right)}$ is indicator function of set $C\left(x_{n}\right)$, and define
$\sigma_{n}(x)=\left\{\begin{array}[]{ll}2R,&x\in\left\{x\mid\rho_{n}(x)\geq 2R\right\}\cap C\left(x_{n}\right)\\ \rho_{n}(x),&x\in\left\{x\mid R<\rho_{n}(x)<2R\right\}\cap C\left(x_{n}\right)\\ \rho_{n}(x)+\frac{\int_{\left\{x\mid\rho_{n}(x)\geq 2R\right\}}\left(\rho_{n}-2R\right)}{\mu\left(\left\{x\mid\rho_{n}(x)\leq R\right\}\cap C\left(x_{n}\right)\right)},&x\in\left\{x\mid\rho_{n}(x)\leq R\right\}\cap C\left(x_{n}\right)\\ 0,&x\in\mathbb{R}^{3}\backslash C\left(x_{n}\right)\end{array}\right$.
Since $\int_{\{x\mid\rho(x)>R\}}R<\int_{\{x\mid\rho(x)>R\}}\rho<\frac{V}{4}$, we know $\mu(\{x\mid\rho(x)>R\})<\frac{V}{4R}<\frac{V}{4}$, then
$\mu\left(\left\{x\mid\rho_{n}(x)\leq R\right\}\cap C\left(x_{n}\right)\right)=\mu\left(C\left(x_{n}\right)\right)-\mu\left(\left\{x\mid\rho_{n}(x)>R\right\}\cap C\left(x_{n}\right)\right)>\frac{3V}{4}$
On the other hand, we have
$\int_{\left\{x\mid\rho_{n}(x)\geq 2R\right\}}\left(\rho_{n}-2R\right)\leq\int_{\left\{x\mid\rho_{n}(x)\geq 2R\right\}}\rho_{n}\leq\int_{\{x\mid\rho(x)>R\}}\rho<\frac{V}{4}$
We obtain
$\frac{\int_{\left\{x\mid\rho_{n}(x)\geq 2R\right\}}\left(\rho_{n}-2R\right)}{\mu\left(\left\{x\mid\rho_{n}(x)\leq R\right\}\cap C\left(x_{n}\right)\right)}<\frac{1}{3}<\frac{3}{4}$
Therefore, if $x\in\left\{x\mid\rho_{n}(x)\leq R\right\}\cap C\left(x_{n}\right)$, we get $\sigma_{n}(x)\leq\rho_{n}(x)+\frac{3}{4}\leq R+\frac{3}{4}<2R$, then we have $\sigma_{n}(x)\leq 2R$ for all $x\in\mathbb{R}^{3}$. One can also check $\int\rho_{n}=\int\sigma_{n}$ (notice $\left\{x\mid\rho_{n}(x)\geq 2R\right\}=\left\{x\mid\rho_{n}(x)\geq 2R\right\}\cap C\left(x_{n}\right)$). Since both $\rho_{n}$ and $\sigma_{n}$ are supported in $C\left(x_{n}\right)$ and $C\left(x_{n}\right)$ has diameter $\frac{\epsilon}{2}$, then diameter of $\operatorname{spt}\left(\rho_{n}-\sigma_{n}\right)$ is not larger than $\frac{\epsilon}{2}$. Thanks to Property ($i$), if $\int\rho_{n}=\int\sigma_{n}>0$, we have $W^{\infty}\left(\rho_{n},\sigma_{n}\right)\leq\frac{\epsilon}{2}$, where $W^{\infty}$ is the Wasserstein $L^{\infty}$ distance on $\mathcal{M}(\mathbb{R}^{3})$.
Let $\sigma=\sum_{n=1}^{\infty}\sigma_{n}$, then $\sigma\in L^{\infty}\left(\mathbb{R}^{3}\right)$ and $\|\sigma\|_{L^{\infty}}\leq 2R$. We know $\int\sigma=\sum_{n}\int\sigma_{n}=\sum_{n}\int\rho_{n}={\int\rho}-{\int_{D}\rho}\stackrel{{\scriptstyle\mu\left(D\right)=0}}{{=}}\int\rho$. In particular, $\sigma$ has the same finite measure as $\rho$, and we can consider the Wasserstein $L^{\infty}$ metric between them.
Claim: $W^{\infty}(\rho,\sigma)<\epsilon$.
In fact, suppose $\int\rho_{n}=\int\sigma_{n}>0$, since $W^{\infty}\left(\rho_{n},\sigma_{n}\right)\leq\frac{\epsilon}{2}$, we can set $\left({S,\Sigma,\nu}\right)=\left({C\left(x_{n}\right),M_{n},c_{n}\mu_{n}}\right)$, where $M_{n}$ is the collection of all Lebesgue measurable sets in $C\left(x_{n}\right)$, $c_{n}$ satisfies $c_{n}\mu_{n}\left({C\left(x_{n}\right)}\right)=c_{n}V=\int\rho_{n}=\int\sigma_{n}>0$, and $\mu_{n}$ satisfies $\mu_{n}(E)=\mu(E\cap C(x_{n}))$ for all measurable sets $E$ in $\mathbb{R}^{3}$, recall $\mu$ is the Lebesgue measure on $\mathbb{R}^{3}$. In particular, we know $c_{n}>0$, and $\mu_{n}(C(x_{n})^{\mathsf{c}})=0$, here $C(x_{n})^{\mathsf{c}}=\mathbb{R}^{3}\setminus C(x_{n})$ is the complement of $C(x_{n})$. Easy to verify that $\mu_{n}$ is a measure in $\mathbb{R}^{3}$ (see for example [17, Exercise 10 in Section 1.3]). Moreover, its restriction to $C(x_{n})$ coincides with the Lebesgue measure on $C(x_{n})$. Therefore, $\left({C\left(x_{n}\right),M_{n},c_{n}\mu_{n}}\right)$ is a (finite) measure space, with $c_{n}u_{n}\left({C\left(x_{n}\right)}\right)=\rho_{n}\left(\mathbb{R}^{3}\right)={\int\rho_{n}}$. Thanks to Remark 4.3, we can choose $\left({C\left(x_{n}\right),M_{n},c_{n}\mu_{n}}\right)$ as the measure space in the definition of $W^{\infty}\left(\rho_{n},\sigma_{n}\right)$ [^1] The reason for choosing such a measure space is to find transport maps $f_{n}$ and $g_{n}$ defined on $C(x_{n})$, which will facilitate the subsequent definition of $\widetilde{f_{n}}$, $\widetilde{g_{n}}$, $f$ and $g$, and the analysis of their properties., then we can find $f_{n}:C\left(x_{n}\right)\rightarrow\mathbb{R}^{3}$ and $g_{n}:C\left(x_{n}\right)\rightarrow\mathbb{R}^{3}$ with ${f_{n}}_{\#}(c_{n}\mu_{n})=\rho_{n},{g_{n}}_{\#}(c_{n}\mu_{n})=\sigma_{n}$, and
$\left\|f_{n}-g_{n}\right\|_{L^{\infty}\left(C\left(x_{n}\right)\right),c_{n}\mu_{n}}<W^{\infty}\left(\rho_{n},\sigma_{n}\right)+\frac{\epsilon}{4}<\frac{3\epsilon}{4}$
Notice since $c_{n}>0$, we have $c_{n}\mu_{n}\ll\mu_{n},\mu_{n}\ll c_{n}\mu_{n}$, and then
$\left\|f_{n}-g_{n}\right\|_{L^{\infty}\left(C\left(x_{n}\right)\right),\mu_{n}}=\left\|f_{n}-g_{n}\right\|_{L^{\infty}\left(C\left(x_{n}\right)\right),c_{n}\mu_{n}}<\frac{3\epsilon}{4}$
Therefore, given a measurable set $B\subset\mathbb{R}^{3}$, we have $c_{n}\mu_{n}\left({f_{n}^{-1}(B)}\right)=\rho_{n}(B)$. Since we know $\rho_{n}$ is supported in $C\left(x_{n}\right)$, then $\rho_{n}(B)=\rho_{n}\left({B\cap C\left(x_{n}\right)}\right)$. Therefore, $c_{n}\mu_{n}\left({f_{n}^{-1}(B)}\right)=\rho_{n}(B)=\rho_{n}\left({B\cap C\left(x_{n}\right)}\right)=c_{n}\mu_{n}\left({f_{n}^{-1}\left({B\cap\left({C\left(x_{n}\right)}\right)}\right)}\right)$. In particular, we know $c_{n}\mu_{n}\left({f_{n}^{-1}((C(x_{n}))^{\mathsf{c}})}\right)=c_{n}\mu_{n}\left({f_{n}^{-1}\left(\emptyset\right)}\right)=c_{n}\mu_{n}\left({\left(\emptyset\right)}\right)=0$, and then $\mu_{n}\left({f_{n}^{-1}((C(x_{n}))^{\mathsf{c}})}\right)=0$ since $c_{n}>0$. Then we can define $\widetilde{f_{n}}:C(x_{n})\rightarrow\mathbb{R}^{3}$ as the following:
$\widetilde{f_{n}}(x)=\left\{\begin{matrix}{f_{n}(x),~~~\text{if}~~f_{n}(x)\in C\left(x_{n}\right)}\\ {0,~~~\text{if}~~f_{n}(x)\in(C\left(x_{n}\right))^{\mathsf{c}}}\end{matrix}\right$.
Notice by construction we know $0\in D$, so $0\notin C(x_{n})$ for any $n$. If $0\notin B$, we can check ${\widetilde{f_{n}}}^{-1}(B)=f_{n}^{-1}\left({B\cap C\left(x_{n}\right)}\right)$, then we have $c_{n}\mu_{n}\left({{\widetilde{f_{n}}}^{-1}(B)}\right)=c_{n}\mu_{n}\left({f_{n}^{-1}\left({B\cap C\left(x_{n}\right)}\right)}\right)=\rho_{n}(B)$. If $0\in B$, we can check ${\widetilde{f_{n}}}^{-1}(B)=f_{n}^{-1}\left({B\cap C\left(x_{n}\right)}\right)\cup f_{n}^{-1}\left({C\left(x_{n}\right)^{\mathsf{c}}}\right)$, and
$\quad\ c_{n}\mu_{n}\left({f_{n}^{-1}\left({B\cap C\left(x_{n}\right)}\right)}\right)$
$\leq c_{n}\mu_{n}\left({{\widetilde{f_{n}}}^{-1}(B)}\right)$
$=c_{n}\mu_{n}\left({f_{n}^{-1}\left({B\cap C\left(x_{n}\right)}\right)\cup f_{n}^{-1}\left({C\left(x_{n}\right)^{\mathsf{c}}}\right)}\right)$
$\leq c_{n}\mu_{n}\left({f_{n}^{-1}\left({B\cap C\left(x_{n}\right)}\right)}\right)+c_{n}\mu_{n}\left({f_{n}^{-1}\left({C\left(x_{n}\right)^{\mathsf{c}}}\right)}\right)$
$=c_{n}\mu_{n}\left({f_{n}^{-1}\left({B\cap C\left(x_{n}\right)}\right)}\right)$
$=\rho_{n}(B)$
Therefore, we know $c_{n}\mu_{n}\left({{\widetilde{f_{n}}}^{-1}(B)}\right)=\rho_{n}(B)=c_{n}\mu_{n}\left({f_{n}^{-1}\left({B\cap C\left(x_{n}\right)}\right)}\right)$ for any measurable set $B\subset\mathbb{R}^{3}$, and thus $\widetilde{f_{n}}_{\#}(c_{n}\mu_{n})=\rho_{n}$. Similarly, we can construct $\widetilde{g_{n}}$ such that $\widetilde{g_{n}}_{\#}(c_{n}\mu_{n})=\sigma_{n}$. Since $\left\|f_{n}-g_{n}\right\|_{L^{\infty}\left(C\left(x_{n}\right)\right),\mu_{n}}<\frac{3\epsilon}{4}$, we can find $U_{n}\subset C\left(x_{n}\right)$ with $\mu_{n}\left(U_{n}\right)=0$, and
${\sup\limits_{C{(x_{n})}\backslash U_{n}}\left|f_{n}-g_{n}\right|}=\left\|{f_{n}-g_{n}}\right\|_{L^{\infty}{({C{(x_{n})}})},\mu_{n}}<\frac{3\epsilon}{4}$
Take $\widetilde{U_{n}}\coloneq U_{n}\cup f_{n}^{-1}\left(\left({C\left(x_{n}\right)}\right)^{\mathsf{c}}\right)\cup g_{n}^{-1}\left(\left({C\left(x_{n}\right)}\right)^{\mathsf{c}}\right)$, since we know $\mu_{n}\left({f_{n}^{-1}((C(x_{n}))^{\mathsf{c}})}\right)=\mu_{n}\left({g_{n}^{-1}((C(x_{n}))^{\mathsf{c}})}\right)=0$, then $\mu_{n}\left(\widetilde{U_{n}}\right)=0$, and we have
[…]
Therefore, $\left\|\widetilde{f_{n}}-\widetilde{g_{n}}\right\|_{L^{\infty}\left(C(x_{n})\right),\mu_{n}}<\frac{3\epsilon}{4}$.
We define $\nu={\sum\limits_{n}{c_{n}\mu_{n}}}$, where $c_{n}$ satisfies $c_{n}\mu_{n}\left({C\left(x_{n}\right)}\right)=c_{n}V=\int\rho_{n}=\int\sigma_{n}$. Notice here we allow $c_{n}=0$. We can see $\nu$ is a measure on $\mathbb{R}^{3}$ and $\nu\left(\mathbb{R}^{3}\right)=\sum_{n}c_{n}\mu_{n}\left(\mathbb{R}^{3}\right)=\sum_{n}c_{n}\mu_{n}\left({C\left(x_{n}\right)}\right)=\sum_{n}\int\rho_{n}=\int\rho=\int\sigma=1<\infty$, thus $\nu$ is a finite measure.
Denote $N_{0}=\left\{n\middle|c_{n}=0\right\}$, and $N_{1}=\left\{n\middle|c_{n}>0\right\}$. We further define $f:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ as
$f(x)=\left\{\begin{array}[]{ll}\widetilde{f_{n}}(x),&x\in C\left(x_{n}\right)\text{with}~n\in N_{1}\\ x,&\text{otherwise}\end{array}\right$.
Similarly, let $g:\mathbb{R}^{3}\rightarrow\mathbb{R}^{3}$ be
$g(x)=\left\{\begin{array}[]{ll}\widetilde{g_{n}}(x),&x\in C\left(x_{n}\right)\text{with}~n\in N_{1}\\ x,&\text{otherwise}\end{array}\right$.
Then $\|f-g\|_{L^{\infty}\left(\mathbb{R}^{3}\right),\nu}\leq\sup\limits_{\{n|c_{n}>0\}}\left\|\widetilde{f_{n}}-\widetilde{g_{n}}\right\|_{L^{\infty}\left(C\left(x_{n}\right)\right),\mu_{n}}\leq\frac{3\epsilon}{4}$. Given a measurable set $B\subset\mathbb{R}^{3}$, let $B_{n}=B\cap C\left(x_{n}\right)$ be the measurable set in $C\left(x_{n}\right),B_{D}=B\cap D$, we know $\mu(D)=0$ implies $\mu\left(B_{D}\right)=0$. We also have if $n\in N_{0}$, then $\rho\left(B_{n}\right)=\rho_{n}\left(B_{n}\right)\leq\rho_{n}\left({C\left(x_{n}\right)}\right)={\int\rho_{n}}=0$, and then
$\rho(B)$ $=\int_{B}\rho=\int_{\cup_{n=1}^{\infty}B_{n}}\rho+\int_{B_{D}}\rho\stackrel{{\scriptstyle\mu\left(B_{D}\right)=0}}{{=}}\int_{\cup_{n=1}^{\infty}B_{n}}\rho$
$={\int_{\cup_{n\in N_{1}}B_{n}}\rho}={\sum_{n\in N_{1}}{\rho\left(B_{n}\right)}}={\sum_{n\in N_{1}}{\rho_{n}\left(B_{n}\right)}}$
$\stackrel{{\scriptstyle\widetilde{f_{n}}_{\#}(c_{n}\mu_{n})=\rho_{n}}}{{=}}{\sum_{n\in N_{1}}{c_{n}\mu_{n}\left({{\widetilde{f_{n}}}^{-1}\left(B_{n}\right)}\right)}}$
We can check $f^{-1}\left(\left\{0\right\}\right)={\bigcup\limits_{n\in N_{1}}{{\widetilde{f}}_{n}}^{-1}}\left(\left\{0\right\}\right)\cup\left\{0\right\}$. Then if $n\in N_{1}$, since ${\widetilde{f_{n}}}_{\#}\left({c_{n}\mu_{n}}\right)=\rho_{n}$, we have $c_{n}\mu_{n}\left({f^{-1}\left(\left\{0\right\}\right)}\right)=c_{n}\mu_{n}\left({{\widetilde{f_{n}}}^{-1}\left(\left\{0\right\}\right)}\right)=\rho_{n}\left(\left\{0\right\}\right)=0$. Easy to check if $n\in N_{0}$, $f^{-1}\left(\left\{0\right\}\right)\cap C\left(x_{n}\right)=\varnothing$, then $\mu_{n}\left({f^{-1}\left(\left\{0\right\}\right)}\right)=0$. Thus we know $v\left({f^{-1}\left\{0\right\}}\right)={\sum_{n}{c_{n}\mu_{n}\left({f^{-1}\left(\left\{0\right\}\right)}\right)}}=0$ (alternatively, this can also be seen by observing that $c_{n}=0$ for all $n\in N_{0}$).
If $n\in N_{1}$, we can check
  1. 1.
    ${\widetilde{f_{n}}}^{-1}\left(B_{n}\right)=f^{-1}\left(B_{n}\right)$
  2. 2.
[…]
  3. 3.
[…]
Therefore, we can compute
$\rho(B)$ $={\sum_{n\in N_{1}}{c_{n}\mu_{n}\left({{\widetilde{f_{n}}}^{-1}\left(B_{n}\right)}\right)}}$
$={\sum_{n\in N_{1}}{c_{n}\mu_{n}\left({f^{-1}\left(B\right)}\right)}}$
$\stackrel{{\scriptstyle c_{n}=0\text{ when }n\in N_{0}}}{{=}}{\sum_{n\in N_{1}}{c_{n}\mu_{n}\left({f^{-1}\left(B\right)}\right)}}+{\sum_{n\in N_{0}}{c_{n}\mu_{n}\left({f^{-1}\left(B\right)}\right)}}$
$={\sum_{n=1}^{\infty}{c_{n}\mu_{n}\left({f^{-1}\left(B\right)}\right)}}$
$=\nu(f^{-1}(B))$
$=f_{\#}\nu(B)$
Similarly, $g_{\#}\nu(B)=\sigma(B)$. Recall $\sigma\in L^{\infty}(\mathbb{R}^{3})$. Therefore, $f_{\#}\nu=\rho$ and $g_{\#}\nu=\sigma$. Moreover, $W^{\infty}(\rho,\sigma)\leq\|f-g\|_{L^{\infty}\left(\mathbb{R}^{3}\right),\nu}\leq\frac{3\epsilon}{4}<\epsilon$. By the construction of $\sigma$ we know if spt $\rho$ is bounded, then $\operatorname{spt}\sigma$ is also bounded. ∎
Remark 4.8.
Lemma 4.7 (vi) facilitates showing that the local minimal energy $E_{J}(\rho)$ of the minimizer $\rho$ is always finite, and then taking derivatives makes sense. We will discuss more about it in Proposition 5.7.
In Theorem 2.25 we assume $(\rho,v)$ minimizes $E(\rho,v)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)\times{V}\left(\mathbb{R}^{3}\right)$, which implies $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)$ by Corollary 2.18. Some might doubt whether this condition is too weak to arrive at the conclusion Theorem 2.25 ($iv$) through (generalized) Lagrange multiplier theorem (Theorem A.4 and Remark A.5) if someone tries to follow the arguments of Auchmuty and Beals [5], where the centers of mass of admissible perturbations are not necessarily 0. But, in fact, it turns out that a local energy minimizer $\rho$ on ${R}_{0}\left(\mathbb{R}^{3}\right)$ is also stable under perturbations which shift its center of mass:
Corollary 4.9 (Center-free and Support-free Minimizer[31, Corollary 5.2]).
If $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)$, then it minimizes $E_{J}(\rho)$ locally on ${R}\left(\mathbb{R}^{3}\right)$.
Proof.
There exists $\delta>0$ such that $E_{J}(\rho)\leq E_{J}(\kappa)$ whenever $\kappa\in{R}_{0}\left(\mathbb{R}^{3}\right)$ with $W^{\infty}(\rho,\kappa)<2\delta$. Now, suppose $\kappa\in{R}\left(\mathbb{R}^{3}\right)$ with $W^{\infty}(\rho,\kappa)<\frac{\delta}{2}$. Thanks to Lemma 4.7 $(ii)$, we know $spt~\kappa$ is also bounded. Lemma 4.7 $(iv)$ shows that $|x(\kappa)|<\frac{\delta}{2}$. Let $\widetilde{\kappa}(x)=\kappa(x+x(\kappa))$, then $\widetilde{\kappa}\in{R}_{0}\left(\mathbb{R}^{3}\right)$ and Lemma 4.7 $(iii)$ shows that $W^{\infty}(\widetilde{\kappa},\kappa)<\frac{\delta}{2}$. By triangle inequality (recall $W^{\infty}$ is a metric), we have $W^{\infty}(\widetilde{\kappa},\rho)\geq W^{\infty}(\widetilde{\kappa},\kappa)+W^{\infty}(\kappa,\rho)<\delta$, i.e. $\widetilde{\kappa}$ lies within $\delta$-neighborhood of $\rho$ in ${R}_{0}\left(\mathbb{R}^{3}\right)$. By translation invariance, $E_{J}(\rho)\leq E_{J}(\widetilde{\kappa})=E_{J}(\kappa)$. ∎
We should point out: suppose $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}_{0}\left(\mathbb{R}^{3}\right)$ and let $\sigma\in L^{\infty}\left(\mathbb{R}^{3}\right)$. Even if the perturbation satisfies $\rho+t\sigma\in{R}\left(\mathbb{R}^{3}\right)$ for $t\in[0,1]$, it may not be $W^{\infty}$-continuous as function of $t$, and we do not know if $E_{J}(\rho+t\sigma)\geq E_{J}(\rho)$ for $t\in[0,1]$; nevertheless, when $\sigma$ is supported on a set which diameter is small enough, with $\int\sigma=0$, then $W^{\infty}(\rho,\rho+t\sigma)$ is small enough for all $t$ thanks to Lemma 4.7 $(i)$, hence $\rho$ minimizes $E_{J}(\rho+t\sigma)$ for all $t$. Therefore, $\sigma$ will then be a useful variation of $E_{J}(\rho)$.
To be precise, we define the set of admissible perturbations, which depend on $\rho$, as:
$P_{\infty}(\rho):=\bigcup_{R<\infty}P_{R}(\rho)$
Here, $P_{R}(\rho)$ is defined as:
[…]
Note that this definition does not a priori require the support of $\sigma$ to have small diameter; such a restriction will be imposed later when needed. The present formulation is convenient because certain papers that do not rely on the $W^{\infty}$ distance do not necessarily require such diameter condition.
One can see $P_{\infty}(\rho)$ is a convex cone.
Remark 4.10.
Given a set $E$ with positive measure, when $R$ is large enough, we can always find a non-zero function $\sigma\in P_{R}(\rho)$, with support in $E$. In fact, since $\rho$ is integrable implies $\rho$ is finite almost everywhere, we have
$1$ $=\int\rho$
$=\sum_{N=0}^{\infty}\int_{\{x\mid N\leq\rho(x)<N+1\}}\rho$
$\geq\sum_{N=0}^{\infty}\int_{\{x\mid N\leq\rho(x)<N+1\}}N$
$=\sum_{N=0}^{\infty}N\mu(\{x\mid N\leq\rho(x)<N+1\})$
$\geq\sum_{N=1}^{\infty}\mu(\{x\mid N\leq\rho(x)<N+1\})$
Notice when $R\rightarrow\infty$, $0\leq\mu\left(\{x\in E\mid\rho(x)>R\}\leq\mu(\{x\in\mathbb{R}^{3}\mid\rho(x)>R\right\}$ $\leq\sum_{N=[R]}^{\infty}\mu(\{x\mid N\leq\rho(x)<N+1\})\rightarrow 0$, hence $\mu(\{x\in E\mid\rho(x)\leq R\})=\mu(E)-\mu(\{x\in E\mid\rho(x)>R\}\rightarrow\mu(E)$. By similar arguments in the following paragraph about how to get $\widetilde{E}$ from $\widetilde{E}_{k}$, we can show when $R$ is large enough, $\mu(\{x\in E\mid\rho(x)\leq R\})>\frac{1}{2}\mu(E)>0$.
Pick one of such $R$, and let $\widetilde{E}_{k}=\{x\in E\mid\rho(x)\leq R,|x|\leq k\}$, then $\left\{\widetilde{E}_{k}\right\}_{k=1}^{\infty}$ is an ascending collection of measurable sets, and $\cup_{k=1}^{\infty}\widetilde{E}_{k}=\{x\in E\mid\rho(x)\leq R\}$, therefore, $\mu(\{x\in E\mid\rho(x)\leq R\})=\lim\limits_{k\rightarrow\infty}\mu\left(\widetilde{E}_{k}\right)$. (see [34, Section 2.5]). When $k$ is large enough, $\mu(\{x\in E\mid\rho(x)\leq R,|x|\leq k\})>\frac{1}{2}\mu(E)>0$, then replace $R$ by another larger $R$ if needed, we have $\mu(\{x\in E\mid\rho(x)\leq R,|x|\leq R\})>0$. We set $\widetilde{E}=\{x\in E\mid\rho(x)\leq R,|x|\leq R\}$, then $\sigma(x)=\left\{\begin{array}[]{ll}1,&x\in\widetilde{E}\\ 0,&\text{otherwise }\end{array}\right$. has the properties we say above. In particular, one can see $0\in P_{\infty}(\rho)$, but $P_{\infty}(\rho)\neq\{0\}$. Notice in $\widetilde{E}$, $\rho(x)$ is bounded and $0\leq\rho(x)\leq R$.
Remark 4.11.
Similar as above, given a set $E$ with positive measure, and $\rho$ has positive mass in $E$ which implies $\mu\{x\in E\mid\rho(x)>0\}>0$, then when $R$ is large enough, let $E_{R}=\{x\in E\mid|x|\leq R,R^{-1}<\rho(x)<R\}$, we have $\mu\left(E_{R}\right)>0$, and $\mathbf{1}_{E_{R}}\in P_{R}(\rho)\subset P_{0}(\rho)$.
Remark 4.12.
Given $\rho\in R(\mathbb{R}^{3})$, and $\sigma\in P_{\infty}(\rho)$, then by definition of $P_{\infty}(\rho)$, we know $\exists R>0$, such that $\sigma\in P_{R}(\rho)$, one can check when t is small enough, $\rho+t\sigma\geq 0$, thus we can also understand from this perspective, that $\rho+t\sigma\geq 0$ is a reasonable perturbation.
5 Analysis of Lagrange Multiplier and Properties of Local Minimizers
This section is divided into three subsections. In the first subsection, we compute the variational derivative of $E_{J}$, and prove the finite energy property of local minimizers, which ensures variational derivative makes sense. In the second subsection, we revisit and modify the proofs of “weak” versions of Theorem 2.25 (iviv-$v$). Here “weak” means (EL) holds true for almost all points, rather than pointwise, in the region we consider. In the third subsection, we point out the particularity of the topology induced by $W^{\infty}$ metric by discussing minimizers’ existence and nonexistence.
5.1 Variational derivative of $E_{J}$ and finite energy of local minimizer
In order to consider the variational derivative $E_{J}^{\prime}(\rho)$ of the energy $E_{J}(\rho)$, we first clarify the space where we want to take derivatives: $X=L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, $U=\{\rho\in X\mid\rho\geq 0,U(\rho)<\left.\infty,G(\rho,\rho)<\infty,T_{J}(\rho)<\infty\right\}$, $W=\left\{\rho\in U\mid\int\rho=1\right\}$.
Remark 5.1.
Suppose $\rho\in X$ is non-negative and positive mass, then by Remark 2.9 we know $I(\rho)>0$, thus $T_{J}(\rho)$ is finite. Thanks to Proposition 2.10 we have $G(\rho,\rho)<\infty$. Therefore,
$\left.\rho\in U\Longleftrightarrow U(\rho)<\infty\Longleftrightarrow E_{J}(\rho)<\infty\right$.
It is helpful if we have a criterion for determining when to obtain $U(\rho)<\infty$.
Lemma 5.2 (Finite Internal Energy).
If a nonnegative function $\sigma$ is in $L^{1}(\mathbb{R}^{3})\cap L^{\infty}(\mathbb{R}^{3})$, then $U(\sigma)<\infty$.
Proof.
Thanks to Remark 2.3, we know $A(s)$ also satisfies (F2) and (F3). In particular, $\exists\delta>0$, such that if $s\in[0,\delta)$, then $A(s)<s^{\frac{4}{3}}$. We know
$U(\sigma)=\int A(\sigma(x))d^{3}x=\int_{\{\sigma<\delta\}}A(\sigma(x))d^{3}x+\int_{\{\delta\leq\sigma\leq C\}}A(\sigma(x))d^{3}x=:U_{1}+U_{2}$
where $C=\|\sigma\|_{L^{\infty}(\mathbb{R}^{3})}$. By the choice of $\delta$ we know $U_{1}\leq\int_{\{\sigma<\delta\}}\sigma^{\frac{4}{3}}(x)d^{3}x\leq\int\sigma^{\frac{4}{3}}(x)d^{3}x$, Since $\sigma\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$, we know by Interpolation Inequality Proposition A.6 that $U_{1}\leq\int\sigma^{\frac{4}{3}}(x)d^{3}x<\infty$. Since $A(s)$ is increasing, we know
$U_{2}$ $\leq\int_{\{\delta\leq\sigma\leq C\}}A(C)d^{3}x$
$\leq A(C)\int_{\{\delta\leq\sigma\leq C\}}\frac{\sigma(x)}{\delta}d^{3}x$
$\leq\frac{A(C)}{\delta}\int_{\{\delta\leq\sigma\leq C\}}{\sigma(x)}d^{3}x$
$\leq\frac{A(C)}{\delta}\|\sigma\|_{L^{1}(\mathbb{R}^{3})}$
$\leq\frac{A(C)}{\delta}<\infty$
Therefore, we know $U(\sigma)<\infty$. ∎
To discuss the differentiability of $E_{J}$, we modify the arguments of Auchmuty and Beals in [5, Section 4] and have the following lemma:
Lemma 5.3 (Differentiability of Energy $E_{J}(\rho)$).
Given $\rho\in W$, we have $P_{\infty}(\rho)\subset P(\rho)$, and $E_{J}(\rho)$ is $P_{\infty}(\rho)$-differentiable at $\rho$. Here $P(\rho)$ and $P_{\infty}(\rho)$-differentiability (i.e. differentiability at $\rho$ in the direction of $P_{\infty}(\rho)$) are provided in Appendix A.1. Moreover, the derivative at $\rho$ is $E_{J}^{\prime}(\rho)$ in the sense that $\forall\sigma\in P_{\infty}(\rho)$, $E_{J}^{\prime}(\rho)(\sigma)=\int E_{J}^{\prime}(\rho)\sigma$ [^3] To remain consistent with the notation in [5, Section 4], we use $E_{J}^{\prime}$ as the symbol for both linear functional and function, provided it does not cause confusion., where on the right hand side the function $E_{J}^{\prime}(\rho)$ is given by
$E_{J}^{\prime}(\rho)(x):=A^{\prime}(\rho(x))-V_{\rho}(x)-\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))$ (28)
Proof.
Given $\sigma\in P_{\infty}(\rho)$, then $\exists R>0$ such that $\sigma\in P_{R}(\rho)$. By construction we know $\sigma\in L^{\infty}\left(\mathbb{R}^{3}\right)$ with compact support, thus by Hölder’s inequality we know $\sigma\in L^{p}\left(\mathbb{R}^{3}\right)$ for $p>0$. Since $\rho\in W$, $\int\rho=1$, then when $t$ is small enough, we have $\int\rho+t\sigma>0$. Thanks to Remark 4.12 and Remark 5.1, we know $\rho+t\sigma$ is nonnegative, $T_{J}(\rho+t\sigma)<\infty$ and $G(\rho+t\sigma,\rho+t\sigma)<\infty$. Moreover, to show $\rho+t\sigma\in U$, it suffices to show $U(\rho+t\sigma)<\infty$.
For almost every $x\in\mathbb{R}^{3},\rho(x)$ is finite, and $\sigma(x)\leq\|\sigma\|_{L^{\infty}}$. For such $x$ fixed and $0<t<1$, by mean value theorem we know $\exists\theta\in(0,t)$, such that
$A(\rho(x)+t\sigma(x))-A(\rho(x))=A^{\prime}(\rho(x)+\theta\sigma(x))t\sigma(x)$ (29)
Set $E_{R}=\left\{x\in\mathbb{R}^{3}\mid\rho(x)\leq R\right$. and $\left.|x|\leq R\right\}$. If $x\notin E_{R}$, by construction $\sigma(x)=0$, thus
$A(\rho(x)+t\sigma(x))-A(\rho(x)=0$
If $x\in E_{R}$, since $A^{\prime}$ is increasing and non-negative, we have
$|A(\rho(x)+t\sigma(x))-A(\rho(x))|\leq A^{\prime}\left(R+\|\sigma\|_{L^{\infty}}\right)\|\sigma\|_{L^{\infty}}\cdot t$
Therefore, we have
$|A\left({\rho+t\sigma}\right)|\leq|A(\rho)|+A^{\prime}\left({R+\left\|\sigma\right\|_{L^{\infty}}}\right)\left\|\sigma\right\|_{L^{\infty}}\cdot\mathbf{1}_{E_{R}}\cdot t$
The right hand side is an integrable function. Then we know $U(\rho+t\sigma)=\int A(\rho+t\sigma)<\infty$. Then we know $\rho+t\sigma\in U$ for $t$ small enough. Therefore, $\sigma\in P(\rho)$ and then $P_{\infty}(\rho)\subset P(\rho)$.
The observations above are also helpful to calculate the derivative. We notice $|\frac{A(\rho(x)+t\sigma(x))-A(\rho(x))}{t}|$ is also bounded by an integrable function $A^{\prime}(R+\|\sigma\|_{L^{\infty}})\|\sigma\|_{L^{\infty}}\cdot\mathbf{1}_{E_{R}}$. Moreover, by (29) we know $\frac{A(\rho(x)+t\sigma(x))-A(\rho(x))}{t}$ converges a.e. to $A^{\prime}(\rho(x))\sigma(x)$ as $t\rightarrow 0$. By dominated convergence theorem, we have
$\lim_{t\rightarrow\infty}\frac{\int A(\rho+t\sigma)-\int A(\rho)}{t}=\int A^{\prime}(\rho)\cdot\sigma.$ (30)
Now consider gravitational interaction energy. Due to Remark 2.11, we know $G(\rho,\rho)=\int\rho V_{\rho}<\infty$ and $G(\rho+t\sigma,\rho+t\sigma)<\infty$, then we compute
$\frac{G(\rho+t\sigma,\rho+t\sigma)-G(\rho,\rho)}{t}=G(\rho,\sigma)+G(\sigma,\rho)-tG(\sigma,\sigma)\rightarrow G(\rho,\sigma)+G(\sigma,\rho)$
Since $G(,\cdot,)$ is a symmetric bilinear form, thus $G(\rho,\sigma)+G(\sigma,\rho)=2G(\sigma,\rho)=2\int\sigma V_{\rho}$.
Consider $T_{J}(\rho)=\frac{J^{2}}{2I(\rho)}$, when $t$ is small enough, we have $\int\rho+t\sigma>0$, thus $I(\rho+t\sigma)>0$. and $T_{J}(\rho+t\sigma)$ is well-defined. Easy to check the mass of $t\sigma$ is $m(t\sigma)=tm(\sigma)$, the center of mass is $\bar{x}(t\sigma)=\bar{x}(\sigma)$, and the moment of inertia is $I(t\sigma)=tI(\sigma)$. Thanks to Lemma 2.15, we can expand $I(\rho+t\sigma)$ and get
$\frac{T_{J}(\rho+t\sigma)-T_{J}(\rho)}{t}$ $=\frac{J^{2}(I(\rho)-I(\rho+t\sigma))}{2I(\rho)I(\rho+t\sigma)\cdot t}$
$=\frac{J^{2}\left(-I(t\sigma)-\left(\frac{m(\rho)m(t\sigma)}{m(\rho)+m(t\sigma)}\right)r^{2}(\bar{x}(\rho)-\bar{x}(t\sigma))\right)}{2I(\rho)I(\rho+t\sigma)\cdot t}$
$=\frac{J^{2}\left(-I(\sigma)-\left(\frac{m(\rho)m(\sigma)}{m(\rho)+m(t\sigma)}\right)r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))\right)}{2I(\rho)I(\rho+t\sigma)}$
$\overset{t\rightarrow 0}{\rightarrow}\frac{J^{2}\left(-I(\sigma)-m(\sigma)\cdot r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))\right)}{2I^{2}(\rho)}$
Consider
$I(\sigma)+m(\sigma)\cdot r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))=$ $\int\sigma\cdot\left(r^{2}(x-\bar{x}(\sigma))+r^{2}(\bar{x}(\rho)-\bar{x}(\sigma))\right)$
$=$ $\int\sigma\cdot(\langle x,x\rangle_{2}-2\langle\bar{x}(\sigma),x\rangle_{2}+\langle\bar{x}(\sigma),\bar{x}(\sigma)\rangle_{2}$
$+\langle\bar{x}(\rho),\bar{x}(\rho)\rangle_{2}-2\langle\bar{x}(\sigma),\bar{x}(\rho)\rangle_{2}+\langle\bar{x}(\sigma),\bar{x}(\sigma)\rangle_{2}$
$-2\langle\bar{x}(\rho),x\rangle_{2}+2\langle\bar{x}(\rho),x\rangle_{2})$
$=$ $\int\sigma\cdot(r^{2}(x-\bar{x}(\rho))+2\langle\bar{x}(\sigma),\bar{x}(\sigma)\rangle_{2}$
$-2\langle\bar{x}(\sigma),x\rangle_{2}-2\langle\bar{x}(\sigma),\bar{x}(\rho)\rangle_{2}+2\langle\bar{x}(\rho),x\rangle_{2})$
$=$ $\int\sigma\cdot r^{2}(x-\bar{x}(\rho))+2m(\sigma)\langle\bar{x}(\sigma),\bar{x}(\sigma)\rangle_{2}-2m(\sigma)\langle\bar{x}(\sigma),\bar{x}(\sigma)\rangle_{2}$
$-2m(\sigma)\langle\bar{x}(\sigma),\bar{x}(\rho)\rangle_{2}+2m(\sigma)\langle\bar{x}(\rho),\bar{x}(\sigma)\rangle_{2}$
$=$ $\int r^{2}(x-\bar{x}(\rho))\cdot\sigma$
Therefore, $\frac{T_{J}(\rho+t\sigma)-T_{J}(\rho)}{t}\rightarrow\int\frac{-J^{2}}{2I^{2}(\rho)}\cdot r^{2}(x-\bar{x}(\rho))\cdot\sigma$ as $t\rightarrow 0$. Collect the results above we finish the proof of Lemma 5.3. ∎
Remark 5.4.
One can generate the results to any $\rho\in U$ with positive mass.
Remark 5.5.
The admissible perturbations are a bit different from $P_{0}(\rho)$ mentioned in [5, Section 4]. There the energy related to momentum is not $T_{J}(\rho)$ but $\frac{1}{2}\int\rho(x)L\left(m_{\rho}(r(x))\right)r^{-2}(x)d^{3}x$, in which the term $r^{-2}(x)$ will be singular when $r(x)\rightarrow 0$. Therefore, the authors require $\sigma(x)=0$ when $r(x)$ is small.
Remark 5.6.
Notice that we rely on the fact that $\sigma$ belongs to $L^{\infty}$ to apply the Dominated Convergence Theorem to obtain (30). Without this assumption, one might consider using Fatou’s Lemma instead, but that would only provide boundedness rather than convergence.
Since an element $\rho\in R(\mathbb{R}^{3})$ might have infinite total energy $E_{J}(\rho)$ — equivalently, infinite internal energy $U(\rho)$ by Remark 5.1 — we have $W\subsetneq R(\mathbb{R}^{3})$. A simple example with infinite total energy could be $\rho(x)\coloneq\rho_{N}(x)=C\frac{1}{|x|^{\alpha}}\mathbf{1}_{\{{{|x|}<N}\}}$ with suitable $C$, $\alpha$, and $N$. Nevertheless, we can show if $\rho$ is a local minimizer, then $\rho\in W$.
Lemma 5.7 (Differentiability at Minimizers).
If $\rho$ minimizes $E_{J}(\rho)$ locally on ${R}\left(\mathbb{R}^{3}\right)$, then $\rho\in W$. In particular, by Lemma 5.3, we know $E_{J}(\rho)$ is $P_{\infty}(\rho)$-differentiable at $\rho$, and the derivative at $\rho$ is $E_{J}^{\prime}(\rho)$ given in (28).
Proof.
Since $\int\rho=1>0$, similar to Remark 5.1 we know $T_{J}(\rho)$ and $G(\rho,\rho)$ are finite. Since $\rho$ is a local minimizer, $\exists\delta>0$, such that $\forall\sigma\in{R}\left(\mathbb{R}^{3}\right)$ with $W^{\infty}(\rho,\sigma)<\delta$, one has $E_{J}(\rho)\leq E_{J}(\sigma)$. Thanks to Lemma 4.7 (vivi), take $\delta$ small enough, we can find a $\sigma\in{R}\left(\mathbb{R}^{3}\right)\cap L^{\infty}\left(\mathbb{R}^{3}\right)$, with $W^{\infty}(\rho,\sigma)<\delta$ and $\int\sigma>0$. Similar as above we have $G(\sigma,\sigma)<\infty,T_{J}(\sigma)<\infty$. Thanks to Lemma 5.2, we know $U(\sigma)<\infty$, and then $E_{J}(\rho)\leq E_{J}(\sigma)<\infty$, which implies $U(\rho)<\infty$. Then we have $\rho\in W$. ∎
5.2 Locally constant Lagrange multiplier
In this subsection, we essentially follow McCann’s arguments [31, Section 5] but with some refinements to prove “weak” versions of Theorem 2.25 Parts (iviv-$v$). The “weak” version serves as the bridge linking the norm of $\rho$ and the norm of potential $V_{\rho}$, which can help to show the global continuity of the $W^{\infty}$ local minimizer $\rho$ as we have seen in the proof of Theorem 2.25
Proposition 5.8 (Locally Constant Lagrange Multiplier [31, Proposition 5.3]).
Let $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$ minimize $E_{J}(\kappa)$ among $\kappa\in{R}\left(\mathbb{R}^{3}\right)$ for which $W^{\infty}(\rho,\kappa)\leq 2\delta$. Let $M$ be an open set with positive measure and with diameter no greater than $2\delta$ which intersects spt $\rho$. There is a unique $\lambda\in\mathbb{R}$ depending on $M$ such that (EL) holds on $M$ a.e.
Proof.
We refer to [31], but slightly modify the arguments and provide a more detailed account (see Remark 5.9). Define the convex cone $P_{loc}:=\left\{\sigma\in P_{\infty}\mid\text{spt }\sigma\subset M\right\}$, which is a cone, and let $U=\{\rho\in X\mid\rho\geq 0,U(\rho)<\left.\infty,G(\rho,\rho)<\infty,T_{J}(\rho)<\infty\right\}$ be the same as in subsection 5.1. Thanks to Lemma 5.3 and Lemma 5.7, we have $\rho\in U,E_{J}(\rho)$ is $P_{loc}$-differentiable at $\rho$. Given $\kappa\in W_{loc}={R}\left(\mathbb{R}^{3}\right)\cap\{\rho+\sigma\in U\mid\sigma\in P_{loc}\}$, where the mass constraint is satisfied, thanks to Lemma 4.7 ($i$) we have $W^{\infty}(\rho,\kappa)\leq\operatorname{diam}M\leq 2\delta$, which shows that $\rho$ minimizes $E_{J}(\kappa)$ on $W_{loc}$. In particular, $\rho$ is a weak local minimizer restricted on $W_{loc}$, which is defined in Appendix A.1. Moreover, since the open set $M$ intersects spt $\rho$, it must carry positive mass under $\rho$, see for example [17, Section 3.4, Exercise 25]. Therefore, similar as Remark 4.11, there is a smaller subset $C\subset M$ of positive measure on which $\rho(x)$ is bounded away from zero and infinity. Let $\mathbf{1}_{C}$ is the indicator function of set $C$, then by construction we have both $\pm\mathbf{1}_{C}\in P_{loc}$. For all $\sigma\in P_{loc}$, let $w=\sigma-\frac{\int\sigma}{\int\mathbf{1}_{C}}\mathbf{1}_{C}$, thanks to Remark 4.12, we know $\rho+tw\in W_{loc}$ when $t$ is small enough. Thanks to Theorem A.4 and Remark A.5, the conditions above imply that there is a unique Lagrange multiplier $\lambda\in R$ such that $\forall\sigma\in P_{loc}$, we have
$\int E_{J}^{\prime}(\rho)\sigma\geq\lambda\int\sigma$ (31)
Claim: (1) $E_{J}^{\prime}(\rho)(x)=A^{\prime}(\rho(x))-V_{\rho}(x)-\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))\geq\lambda$ on $M$ a.e. and (2) $E_{J}^{\prime}(\rho)=\lambda$ on $M\cap\{x\mid\rho(x)>0\}$ a.e.
In fact, If $E_{J}^{\prime}(\rho)<\lambda$ on a subset $K\subset M$ which had positive measure, similar as Remark 4.10, this subset may be taken slightly smaller so that $\rho$ is bounded on $K$. Then we take $\sigma=\mathbf{1}_{K}\in P_{loc}$ and would see it contradict (31), thus Claim (1) holds true. On the other hand, if $E_{J}^{\prime}(\rho)>\lambda$ on a subset $K\subset M$ with positive measure and $\rho>0$ in $K$, then similarly $K$ may be taken slightly smaller so that $\rho$ is bounded away from zero and infinity on $K$; in this case $\mathbf{1}_{K}\in P_{loc}$ can make a contradiction to (31) again, thus Claim (2) holds true.
By Remark 2.5 we know $A^{\prime}(0)=0$, and $A^{\prime}(s)>0$ when $s>0$. If $V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda>0$, then by Claim (1) $A^{\prime}(\rho(x))\geq V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda>0$, which implies $\rho>0$. By Claim (2) we can replace inequality by equation and get $A^{\prime}(\rho(x))=V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda$. If $V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda\leq 0$, then $\rho=0$, otherwise $\rho>0$ implies again $0<A^{\prime}(\rho(x))=V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda\leq 0$, which leads to a contradiction. Therefore, we have (EL), that is $A^{\prime}(\rho(x))=\left[V_{\rho}(x)+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda\right]_{+}$ holds for almost all $x\in M$. Since $M$ has positive measure and $\rho$ in $M$ has positive mass, we know there exists a $x_{0}$ such that $\rho(x_{0})>0$ and thus $A^{\prime}(\rho(x_{0}))>0$, and (EL) holds at $x_{0}$. In particular, $0<A^{\prime}(\rho(x_{0}))=V_{\rho}(x_{0})+\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x_{0}-\bar{x}(\rho))+\lambda$. It is clear that $\lambda$ is unique and given by this expression . ∎
Remark 5.9.
Compared to the McCann’s proof, we have added the following discussions:
1. When deriving inequality (31), McCann simply referenced Theorem A.4. However, the scope of Theorem A.4 does not fully cover the situation here. Therefore, we supplemented the proof with Remark A.5 and utilized both Theorem A.4 and Remark A.5 to reach the conclusion.
2. After deriving Claim (1) and Claim (2), McCann directly asserted Proposition 5.8’s conclusion that (EL) holds on $M$ a.e. In contrast, we added an explanation detailing how this conclusion is obtained.
Remark 5.10.
We recall again $\rho+tw\in{R}\left(\mathbb{R}^{3}\right)$ may not be $W^{\infty}$-continuous as function of $t$. In another word, whether a local minimizer is also a weak local minimizer remains unclear in general. However, at least we know a minimizer restricted to $W_{loc}$ is a weak local minimizer restricted to $W_{loc}$, so that we can apply Theorem A.4 and Remark A.5. This is why we first consider the local case Proposition 5.8 (the diameter of $M$ is small) before addressing the next case Proposition 5.11.
Proposition 5.11 (Componentwise Constant Lagrange Multiplier [31, Proposition 5.4]).
Let $\rho\in{R}_{0}\left(\mathbb{R}^{3}\right)$ minimize $E_{J}(\kappa)$ among $\kappa\in{R}\left(\mathbb{R}^{3}\right)$ for which $W^{\infty}(\rho,\kappa)<2\delta$. $\Omega_{i}$ is one of the connected components of the $\delta$-neighbourhood of spt $\rho$. Then there is a constant $\lambda_{i}<0$ such that (EL) holds a.e. on $\Omega_{i}$.
Proof.
Given $y\in\Omega_{i}$, we know the ball $B_{\delta}(y)$ intersects spt $\rho$. Thus Proposition 5.8 guarantees a unique $\lambda(y)$ such that (EL) holds a.e. on $B_{\delta}(y)$. The claim is that $\lambda(y)$ is independent of $y$. In fact, fix $y_{0}\in\Omega_{i}$, since $B_{\delta}(y_{0})$ is open, it will also be true that a slightly smaller ball $B_{\delta-\epsilon}(y_{0})$ intersects spt $\rho$. If $|x-y_{0}|<\epsilon$, then $M=B_{\delta}(x)\cap B_{\delta}(y_{0})$ intersects spt $\rho$ since $B_{\delta-\epsilon}(y_{0})\subset M$, and then we know $x\in\Omega_{i}$. $M$ has positive measure, thanks to Proposition 5.8, the uniqueness of $\lambda$ corresponding to $M$ forces $\lambda(x)=\lambda(y_{0})$. Thus $\lambda(y)$ is locally constant. As a result, the disjoint sets $C=\left\{x\in\Omega_{i}\mid\lambda(x)=\lambda(y_{0})\right\}$ and $D=\left\{x\in\Omega_{i}\mid\lambda(x)\neq\lambda(y_{0})\right\}$ are both open. Since $\Omega_{i}=C\cup D$ is connected and $y_{0}\in C$ implies $C$ is not empty, $C=\Omega_{i}$. Defining $\lambda_{i}:=\lambda(y_{0})$,(EL) must be satisfied a.e. on $\Omega_{i}$.
Now we show $\lambda<0$ : The distance between any point on the boundary of $\Omega_{i}$ and spt $\rho$ cannot be smaller than $\delta$. Since spt $\rho$ is bounded, $\Omega_{i}$ has non-empty boundary, and it follows that $\rho(x)=0$ on a set of positive measure in $\Omega_{i}$. On the other hand, $A^{\prime}(\rho)\rho-A(\rho)=P(\rho)$ implies $A^{\prime}(\rho(x))$ vanishes only if $\rho(x)=0$. $\lambda\geq 0$ in (EL) would imply $\rho>$ 0 a.e. on $\Omega_{i}$, a contradiction. ∎
Remark 5.12.
Notice we discuss above the connected components of the $\delta$-neighbourhood of spt $\rho$. But the result can be easily converted to the case about $\delta$-neighbourhood of connected components of spt $\rho$.
Remark 5.13.
From the above proof, we can see the power of Wasserstein metric, which can extend the Euler-Lagrange equation (EL) to a $\delta$-neighborhood of connected component of spt $\rho$, which help us to show the global continuity of $\rho$ and then (EP’) holds as we discuss in Theorem 2.25. Moreover, it plays a role in the process of proving the Lagrange multiplier is negative. Actually, there is another way to show Lagrange multipliers should be negative, given by Auchmuty and Beals [5], with different condition.
5.3 Minimizers’ existence and nonexistence
We recall again under the topology induced by $W^{\infty}$ metric, it is not entirely certain whether a local minimizer is a weak local minimizer, see Remark 5.10. For comparison, we prove a local minimizer is a weak local minimizer under the topology inherited from topological vector space. Furthermore, we demonstrate that, in this case, a local minimum either does not exist or is not physically meaningful.
Denote $X=L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)\cap L^{1}\left(\mathbb{R}^{3}\right)$, $\widetilde{U}=\{\rho\in X\mid\rho\geq 0\}$. We first show in the topology inherited from the topological vector space, a local minimizer is a weak local minmizer.
Lemma 5.14.
If the topology on ${R}\left(\mathbb{R}^{3}\right)$ is inherited from a topological vector space, and $\rho$ is a local minimizer for $E_{J}(\rho)$ restricted on ${R}\left(\mathbb{R}^{3}\right)$, then $\rho$ is a weak local minimizer restricted on ${R}\left(\mathbb{R}^{3}\right)$, which is defined in Appendix A.1.
Proof.
We denote $\left.N\left(\mathbb{R}^{3}\right)=\left\{{\sigma\in L^{\frac{4}{3}}\left(\mathbb{R}^{3}\right)}\right|\int\sigma=0\right\}$. We recall the setup in Appendix A.1, that $P(\rho)=\left\{{\sigma\in X}\right|\exists\epsilon(\sigma),~\text{such that}~\rho+t\sigma\in\widetilde{U},~0\leq t<\epsilon(\sigma)\}$.
Since the topology on $R(\mathbb{R}^{3})$ is inherited from a topological vector space, we know there is a topological vector space $Y$ with topology $\mathcal{T}$, such that $R(\mathbb{R}^{3})$ is a subset of $Y$, with the collection $\mathcal{T}_{R}=\left\{R\left(\mathbb{R}^{3}\right)\cap U\middle|U\in\mathcal{T}\right\}$ being the topology on $R(\mathbb{R}^{3})$. (See [33, Chapter 1, Section 16]).
Since $\rho$ is a local minimizer restricted on $R(\mathbb{R}^{3})$, then $\exists U_{R}\in\mathcal{T}_{R}$, such that $\rho\in U_{R}$ and $E_{J}(\rho)\leq E_{J}(\kappa)$ for all $\kappa\in U_{R}$. Notice that $U_{R}=R(\mathbb{R}^{3})\cap U$ for some $U\in\mathcal{T}$, and thus $\rho\in U$. Given $\sigma\in P(\rho)\cap N(\mathbb{R}^{3})$, we know $\rho+t\sigma\in R(\mathbb{R}^{3})\subset Y$. Since $Y$ is a topological vector space, we know when $t$ is small enough, $\rho+t\sigma\in U$. Therefore, $\rho+t\sigma\in R(\mathbb{R}^{3})\cap U=U_{R}$, and then $E_{J}(\rho)\leq E_{J}(\rho+t\sigma)$ for $t$ small enough. Then we know $\rho$ is a weak local minimizer restricted on $R(\mathbb{R}^{3})$. ∎
Proposition 5.15 (No Local Minimizers with Finite Energy in a Vector Space Topology [31, Remark 3.7]).
For $J>0$, if the topology on ${R}\left(\mathbb{R}^{3}\right)$ is inherited from a topological vector space, then there is no local minimizer for $E_{J}$ restricted on ${R}\left(\mathbb{R}^{3}\right)$, with $E_{J}(\rho)<\infty$.
Remark 5.16.
Note that, unlike in McCann’s statement, here we first assume the local minimal energy $E_{J}(\rho)$ is finite, so that we can mimic the arguments in Lemma 5.3 and then make contradiction.
Proof of Proposition 5.15.
Suppose $\rho$ is a local minimizer restricted on $R(\mathbb{R}^{3})$, with $E_{J}(\rho)<\infty$. Thanks to Remark 5.1, we know $U(\rho)<\infty$, that is, $A(\rho)$ is integrable. Therefore, similar arguments to Lemma 5.3, we know $P_{\infty}(\rho)\subset P(\rho)=\{\sigma\in X\mid\exists\epsilon(\sigma),\text{such that~}\rho+t\sigma\in\widetilde{U},0\leq t<\epsilon(\sigma)\}$, and $E_{J}(\rho)$ is $P_{\infty}(\rho)$-differentiable at $\rho$. Thanks to Lemma 5.14, we know $\rho\in{R}\left(\mathbb{R}^{3}\right)$ is a weak local minimizer restricted on $R(\mathbb{R}^{3})$. Let $\sigma\in P_{\infty}(\rho)$, similar to the discussion in Proposition 5.8, we can again derive (31) here. Notice there are enough functions $\sigma$ we obtain (EL)
$A^{\prime}(\rho(x))=\left[\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+V_{\rho}(x)+\lambda\right]_{+}$
is satisfied a.e. on $\mathbb{R}^{3}$ for a fixed $\lambda$. Then we have for almost all $x\in\mathbb{R}^{3}$,
$A^{\prime}(\rho(x))\geq\frac{J^{2}}{2I^{2}(\rho)}r^{2}(x-\bar{x}(\rho))+\lambda$
Since $A^{\prime}$ is strictly increasing, this inequality implies for any $C>0$, $\exists R>0$, such that $\rho(x)>C$ almost everywhere in the region $r(x)>R$ , which contradicts the fact that $\rho\in{R}\left(\mathbb{R}^{3}\right)$ has mass 1. ∎
Remark 5.17.
Observing that A is a linear cone, it is intuitively reasonable to suspect that the minimum of f over A A may not exist. Therefore, we instead consider the minimization of f f over B B.
Remark 5.18.
Inspired by McCann’s Remark [31, Remark 3.7], we can also understand physically Proposition 5.15 as the following: suppose there is a small portion of mass in a rotating star which is moved to a very distant location, that is, from $\kappa=\rho(x)+\sigma(x)$ to $\kappa_{N}=\rho(x)+\sigma(x-N)$ for $N$ large. Then intuitively as $N\rightarrow\infty$, the internal energy $U$ and gravitational interaction energy $G$ will not change to much, while $I(\kappa_{N})\rightarrow\infty$, which implies the kinetic energy $T_{J}\rightarrow 0$, leading to the decrease of $E_{J}$ in the end.
Remark 5.19.
In McCann’s discussion [31, Remark 3.7], the conclusion considers $\rho$ as a local minimizer restricted to $R_{0}(\mathbb{R}^{3})$ rather than $R(\mathbb{R}^{3})$. However, as revealed by the proof above, if we focus solely on the restriction to $R_{0}(\mathbb{R}^{3})$, it becomes challenging to demonstrate the existence of sufficiently many functions $\sigma$ to ensure that the (EL) holds. Consequently, it is difficult to establish the non-existence of such $\rho$.
Remark 5.20.
From the proof above we notice that the issue comes from the fact that even if $\sigma(x)$ remains positive when $x$ is far away from spt $\rho$, $\rho+t\tilde{\sigma}$ is still continuous with respect to $t$, which implies local minimizer $\rho$ is also a weak local minimizer. Fortunately, this situation does not occur if ${R}\left(\mathbb{R}^{3}\right)$ is topologized via the Wasserstein $L^{\infty}$ metric. This is because if $\sigma(x)$ remains positive when $x$ is far away from spt $\rho$, then $\rho+t\tilde{\sigma}$ might not be continuous with respect to t, or $\rho+t\tilde{\sigma}$ might not fall within the neighborhood of $\rho$. In another word, a local minimizer is not necessary to be a weak local minimizer as we have seen in Remark 5.10.
From the proof of Proposition 5.15, we obtain the nonexistence of local minima with finite energy $E_{J}(\rho)$. The finiteness assumption is essential for following the arguments in Lemma 5.3 to compute derivative of $U(\rho)$. However, if the minimal energy $E_{J}(\rho)$ blows up, i.e., becomes infinite, the arguments in Lemma 5.3 can no longer be utilized.
One might hope to adopt an argument similar to Lemma 5.7 to show that if $\rho$ is a local minimum, its energy must necessarily be finite. While this is achievable under the topology induced by the Wasserstein $L^{\infty}$ distance, it remains uncertain whether such a $\rho$ with finite energy can be found under the topology inherited from the topological vector space. The following theorem leads to a conjecture that the statement regarding local minimizer (w.r.t. topology inherited from the topological vector space) having finite energy might not hold.
We first make an additional assumption of $P(\rho)$, which helps us to compare $A(\rho)$ and $A(\lambda\rho)$:
  1. (F5)
    There exists $\lambda\in(0,1)$, such that
    $\limsup\limits_{\rho\rightarrow\infty}\frac{\int_{\lambda\rho}^{\rho}{P(\tau)\tau^{-2}d\tau}}{\int_{0}^{\lambda\rho}{P(\tau)\tau^{-2}d\tau}}$$\infty$ (32)
Remark 5.21.
One can check this condition will be satisfied in the case $P(\rho)=K\rho^{\gamma}$.
Remark 5.22.
Notice that if we can find a $\lambda$ such that (32) holds, then $\forall\widetilde{\lambda}\in(\lambda,1)$, (32) also holds.
Proposition 5.23 (Existence of Weak Local Minimizer when Energy Blows Up).
If $P(\rho)$ in addition satisfies (F5), and $\rho\in R_{0}(\mathbb{R}^{3})$ satisfies $E_{J}(\rho)=\infty$. Then $\forall\sigma\in P(\rho)=\{\sigma\in X\mid\exists\epsilon(\sigma),\text{such that~}\rho+t\sigma\in\widetilde{U},0\leq t<\epsilon(\sigma)\}$, $\exists\delta>0$, such that $\forall t\in(0,\delta)$, we have $E_{J}(\rho+t\sigma)=\infty$. In particular, $\rho$ is a weak local minimizer (see Appendix A.1) for $E_{J}(\rho)$ on $R(\mathbb{R}^{3})$.
Proof.
If $E_{J}(\rho)=\infty$, due to Remark 5.1, we know $U(\rho)={\int{A(\rho)}}=\infty$. Since $\rho$ has compact support, $A(s)$ is an increasing function of $s$, and $\rho=0$ if and only if $A(\rho)=0$, then we know $\forall N>0$,
${\int_{\{{\rho\leq N}\}}{A(\rho)}}\leq A(N){\int_{\{{\rho\leq N}\}}1}\leq A(N){\int_{spt~\rho}1}=A(N)\mu\left({spt~\rho}\right)<\infty$
where $\mu$ is Lebesgue measure.
Since we know ${\int{A(\rho)}}={\int_{\{{\rho\leq N}\}}{A(\rho)}}+{\int_{\{{\rho>N}\}}{A(\rho)}}=\infty$, we know $\int_{\{{\rho>N}\}}{A(\rho)}$$\infty$ for all $N>0$.
On the other hand, suppose $\sigma\in P(\rho)$, then we know $\forall t\in(0,\epsilon(\sigma))$, we have $\rho+t\sigma\geq 0$, then $\sigma\geq-\frac{\rho}{t}\geq-\frac{\rho}{\epsilon(\sigma)}$. If $t\rightarrow 0^{+}$, notice $A(s)$ is a increasing function, we have $U\left({\rho+t\sigma}\right)={\int{A(\rho+t\sigma)}}\geq{\int{A\left({\left({1-\frac{t}{\epsilon(\sigma)}}\right)\rho}\right)}}\geq{\int_{\{{\rho>N}\}}{A\left({\left({1-\frac{t}{\epsilon(\sigma)}}\right)\rho}\right)}}$ for all $N$. Since $t>0$, then $\lambda=1-\frac{t}{\epsilon(\sigma)}<1$, and ${\lim\limits_{t\rightarrow 0^{+}}\lambda}=1$.
Due to (F5), we denote $L\coloneq{\limsup\limits_{\rho\rightarrow\infty}\frac{\int_{\lambda\rho}^{\rho}{P(\tau)\tau^{-2}d\tau}}{\int_{0}^{\lambda\rho}{P(\tau)\tau^{-2}d\tau}}}<\infty$, and claim: $\exists\eta\in(0,1)$ and $N_{1}>0$, such that if $\lambda\in(\eta,1)$ and $\rho>N_{1}$, we have $A(\lambda\rho)\geq\frac{\lambda}{L+1}A(\rho)$ (for the case $P(\rho)=K\rho^{\gamma}$ with $\gamma>\frac{4}{3}$ and hence $A(\rho)=\frac{K}{\gamma-1}\rho^{\gamma}$ (cf. (2)), the inequality can be verified directly).
In fact, given $\lambda<1$, by the definition of $A$ (1), we have
$A\left({\lambda\rho}\right)=\lambda\rho{\int_{0}^{\lambda\rho}{P(\tau)\tau^{-2}d\tau}}=\lambda\rho{\int_{0}^{\rho}{P(\tau)\tau^{-2}d\tau}}-\lambda\rho{\int_{\lambda\rho}^{\rho}{P(\tau)\tau^{-2}d\tau}}=\lambda A(\rho)-\lambda\rho{\int_{\lambda\rho}^{\rho}{P(\tau)\tau^{-2}d\tau}}$
Due to Remark 5.22, we can find $\eta\in(0,1)$ and $N_{1}>0$, such that if $\lambda\in(\eta,1)$ and $\rho>N_{1}$, we have
$\lambda\rho{\int_{\lambda\rho}^{\rho}{P(\tau)\tau^{-2}d\tau}}\leq\lambda L\rho{\int_{0}^{\lambda\rho}{P(\tau)\tau^{-2}d\tau}}=\lambda L\rho{\int_{0}^{\rho}{P(\tau)\tau^{-2}d\tau}}-\lambda L\rho{\int_{\lambda\rho}^{\rho}{P(\tau)\tau^{-2}d\tau}}$
. Then we have
$\lambda\rho{\int_{\lambda\rho}^{\rho}{P(\tau)\tau^{-2}d\tau}}\leq\frac{\lambda L}{L+1}\rho{\int_{0}^{\rho}{P(\tau)\tau^{-2}d\tau}}=\frac{\lambda L}{L+1}A(\rho)$
Therefore, by (1), we obtain the claim:
$A\left({\lambda\rho}\right)=\lambda A(\rho)-\lambda\rho{\int_{\lambda\rho}^{\rho}{P(\tau)\tau^{-2}d\tau}}\geq\frac{\lambda}{L+1}A(\rho)$
Due to this claim, we know $\exists\delta>0$, such that when $t\in(0,\delta)$, we have
$U\left({\rho+t\sigma}\right)\geq{\int_{\{{\rho>N_{1}}\}}{A\left({\left({1-\frac{t}{\epsilon(\sigma)}}\right)\rho}\right)}}\geq\frac{1}{L+1}\left({1-\frac{t}{\epsilon(\sigma)}}\right){\int_{\{{\rho>N_{1}}\}}{A(\rho)}}=\infty$
Thanks to Proposition 2.10, we know $G(\rho+t\sigma,\rho+t\sigma)<\infty$, then we know $E_{J}(\rho+t\sigma)=\infty$ for $t\in(0,\delta)$. ∎
Remark 5.24.
Notice that, in Proposition 5.23, we do not require the topology being inherited from a topological vector space. In fact, when defining a weak local minimizer (see Appendix A.1), we do not utilize any specific properties of any topology. (Although we call it “weak” because, under the concept of a topological vector space, a local minimizer can imply a weak local minimizer. See, for example, Lemma 5.14.)
We can also recall Remark 5.10, which states that, in the topology induced by $W^{\infty}$ distance, a local minimizer is not necessarily a weak local minimizer. Therefore, it is interesting to compare Proposition 5.23 and Lemma 5.7.
Appendix A Appendix
Here we introduce some preliminary knowledge which are used in the paper.
A.1 Calculus of Variations in Vector Spaces
Since the functional we deal with is not differentiable in any usual sense, we begin with a less restrictive notion of differentiability.
Let $X$ be a real vector space, $U$ a subset of $X$, and $E$ an extended real-valued function defined on $U$. Given $u\in U$, let $P(u)=\{v\in X\mid\exists\epsilon(v)>0,\text{ such that }u+tv\in U\text{ for }t\in[0,\epsilon(v))\}$. This set is a cone in $X$. If $P$ is any cone in $X$, let $X(P)$ be the linear subspace of $X$ generated by $P$. We say that $E$ is $P$-differentiable at $u$ if $P\subset P(u)$ and there is a linear functional $\mu$ on $X(P)$ such that $\forall v\in P$
$\lim\limits_{t\rightarrow 0^{+}}t^{-1}\{E(u+tv)-E(u)\}=\mu(v),$
The linearity of $\mu$ is needed in our analysis; its precise domain is not essential here. We denote the “derivative” $\mu$ by $E_{u}^{\prime}$—with $P$ understood from context.
We say $E$ has a weak local minimum at $u\in U$ if for each $v\in P(u)$, there exists $\{t_{n}\}$ such that $t_{n}>0$, ${\lim\limits_{n\rightarrow\infty}t_{n}}=0$ and $E\left({u+t_{n}v}\right)\geq E(u)$ for $n$ large enough. And $u$ is called weak local minimizer
Remark A.1.
If $u$ is a weak local minimizer, then one can check
$\lim_{t\rightarrow 0^{+}}\inf E(u+tv)\geq E(u)$
In fact, the latter corresponds to the definition of weak local minimizer in [5, Section 2]. However, this limited-based definition does not fully align with our intuitive understanding of a “minimizer”. For instance, consider $E(u)=-u^{2}$; according to this limited-based definition, $u=0$ would qualify as a weak local minimizer. Yet, this is not a reasonable conclusion. Our definition, on the other hand, helps exclude such case.
It is only through our definition of weak local minimizer that we can truly establish the following proposition in [5, Section 2]:
Proposition A.2 ( [5, Section 2] ).
Suppose $E$ has a weak local minimum at $u$ and is $P$-differentiable at $u$. Then $E_{u}^{\prime}(v)\geq 0$ for all $v\in P$.
Proof.
The proof can be established using a proof by contradiction, or by referring to Theorem A.4 below, specifically its proof for $E_{u}^{\prime}(w)\geq 0$. ∎
Remark A.3.
Generally speaking, if $X$ is a topological vector space, that is, $u+tv$ is continuous w.r.t. $t$, then one can prove that a local minimizer is a weak local minimizer. For reference, see Lemma 5.14. What’s more, the “weak” nature of a weak local minimizer, compared to a local minimizer, is also reflected in the following:
To check $u$ is weak local minimizer, we just need to show $E(u)$ is a minimum over $E(u+t_{n}v)$ for some $t_{n}$} in any arbitrary direction $v\in P(u)$, rather than for all small $t$ in the direction $v$. Moreover, we do not need to specify the topology or find a neighborhood of $\rho$.
Although it is “weak”, a weak local minimizer retains the possibility of computing or analyzing derivatives, as we can see in Proposition A.2 and Theorem A.4.
Let $W=\left\{u\in U\mid g(u)=M\right\}$, where $g$ is a linear functional on $X$ and $M$ is a constant. Then similarly, we say $E$ has a weak local minimum at $u$ restricted to $W$ if $u\in W$ and for each $v\in P(u)$, there exists $t_{n}$} such that $t_{n}>0$, ${\lim\limits_{n\rightarrow\infty}t_{n}}=0$, and for $n$ large enough, we have $u+t_{n}v\in W$ [^4] Let $N=\left\{u\in U\mid g(u)=0\right\}$, we can check “equivalent to $v\in P(u)$ such that $u+tv\in W$ for $t$ small enough” is $v\in P(u)\cap N$” . However, when considering the extended result in the Remark A.5, this criterion can be different and not equivalent, and it requires a more detailed discussion, similar to what is done in the proof of Proposition 5.8., and $E\left({u+t_{n}w}\right)\geq E(u)$. Furthermore, we have a result for weak local minimum restricted to $W$:
Theorem A.4 ((Generalized) Lagrange Multiplier Theorem [5, Section 2]).
Suppose $E$ is $P$-differentiable at $u\in U$ and that its restriction to $W$ has a weak local minimum at $u$. Suppose also that $P$ is convex and that there is a $u_{0}\in P$ such that $u_{0}\in P$ and $g\left(u_{0}\right)\neq 0$. Then there is a unique constant $\lambda$ such that $E_{u}^{\prime}(v)\geq\lambda g(v)$, all $v\in P$.
Proof.
Given $v\in P$, let $w=v-\frac{g(v)}{g(u_{0})}u_{0}$, then $g(w)=0$. Since $P$ is a cone and $\pm u_{0}\in P$, we know $\frac{g(v)}{g(u_{0})}u_{0}\in P$. Furthermore, since $P$ is convex and $v\in P$, we know $\frac{1}{2}v-\frac{1}{2}\frac{g(v)}{g(u_{0})}u_{0}\in P$. Using again the fact $P$ is a cone, we have $w=v-\frac{g(v)}{g(u_{0})}u_{0}\in P\subset P(u)$.
Therefore, by the definition of $P(u)$, we know $u+tw\in U$ for small positive $t$, thus $u+tw\in W$. $E$ is $P$-differentiable at $u\in U$, then we know $E_{u}^{\prime}(w)$ exists.
Claim: $E_{u}^{\prime}(w)\geq 0$.
In fact, if $E_{u}^{\prime}(w)<0$, by definition of $E_{u}^{\prime}(w)$ we know exists $\delta>0$, such that if $0<t<\delta$, we have
$E\left({u+tw}\right)<E(u)$
On the other hand, $u$ is a weak local minimizer restricted to $W$, then we know $\exists{t_{n}}$ such that $t_{n}>0$, ${\lim\limits_{n\rightarrow\infty}t_{n}}=0$ and $E\left({u+t_{n}w}\right)\geq E(u)$ for $n$ large enough. It leads to a contradiction.
Therefore, $E_{u}^{\prime}(w)\geq 0$. Then we know
$0\leq E_{u}^{\prime}(w)=E_{u}^{\prime}(v)-\frac{g(v)}{g(u_{0})}E_{u}^{\prime}(u_{0})$
Thus we may take $\lambda=\frac{E_{u}^{\prime}(u_{0})}{g(u_{0})}=E_{u}^{\prime}(\frac{u_{0}}{g(u_{0})})$, and get
$E_{u}^{\prime}(v)\geq\lambda g(v)$
On the other hand, if $\lambda$ does satisfy the above, then $E_{u}^{\prime}(u_{0})\geq\lambda g(u_{0})$ and $E_{u}^{\prime}(-u_{0})\geq-\lambda g(u_{0})$. Thus $\lambda$ has to be $E_{u}^{\prime}(\frac{u_{0}}{g(u_{0})})$ and unique. ∎
Remark A.5.
If we go through the proof above, we can find that $W$ can be a subset of $\left\{u\in U\middle|g(u)=M\right\}$. At this point, we need to ensure that $u+t_{n}w$ in the proof above indeed lies within $W$ when $n$ is sufficiently large.
A.2 Properties of Sobolev Spaces
We recall some properties of Sobolev spaces as well as of $L^{p}$ spaces. Some of them will be used frequently in this thesis.
Proposition A.6 (Interpolation Inequality [9, Section 4.2]).
If $f\in L^{p}\cap L^{q}$ with $1\leq p\leq q\leq\infty$, then $f\in L^{r}$ for all $r,p\leq r\leq q$, and the following interpolation inequality holds:
$\|f\|_{r}\leq\|f\|_{p}^{\alpha}\|f\|_{q}^{1-\alpha},\text{ where }\frac{1}{r}=\frac{\alpha}{p}+\frac{1-\alpha}{q},0\leq\alpha\leq 1$
Proposition A.7 (Hardy-Littlewood-Sobolev Inequality [8, Theorem 1.7]).
Let $1<p,r<\infty$ and $0<\alpha<n$ be such that $\frac{1}{p}+\frac{\alpha}{n}=\frac{1}{r}+1$. $\exists C_{p,\alpha,n}>0$, such that
$\left\||\cdot|^{-\alpha}*f\right\|_{L^{r}\left(\mathbb{R}^{n}\right)}\leq C_{p,\alpha,n}\|f\|_{L^{p}\left(\mathbb{R}^{n}\right)}$ (33)
Remark A.8.
Notice that $|x|^{-\alpha}$ is not in $L^{\frac{n}{\alpha}}(\mathbb{R}^{3})$, hence Young’s Inequality [9, Theorem 4.33 and Exercise 4.30] is not applicable. But Hardy-Littlewood-Sobolev Inequality Proposition A.7 can also fail at the case $p=1,r=\infty$, or $\alpha=n$.
Notice that Hardy-Littlewood-Sobolev Inequality Proposition A.7 can fail when $r=\infty$. To estimate the bound in $L^{\infty}$, we introduce the following proposition:
Proposition A.9 (Bound of Potential [5, Proposition 5]).
Suppose $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{p}\left(\mathbb{R}^{3}\right)$. If $1<p\leq\frac{3}{2}$, then $\forall r\in\left(3,\frac{3p}{3-2p}\right),V_{\rho}\in L^{r}\left(\mathbb{R}^{3}\right)$, and $\exists 0<b_{r}<1,0<c_{r}<1,C>0$, such that
$\left\|V_{\rho}\right\|_{L^{r}}\leq C\left(\|\rho\|_{L^{1}}^{b_{r}}\|\rho\|_{L^{p}}^{1-b_{r}}+\|\rho\|_{L^{1}}^{c_{r}}\|\rho\|_{L^{p}}^{1-c_{r}}\right)$ (34)
If $p>\frac{3}{2}$, then $V_{\rho}$ is bounded and continuous and satisfies (34) with $r=\infty$.
Proof.
One can see the detailed proof for the case $1<p\leq\frac{3}{2}$ in [5]. For the case $p>$ $\frac{3}{2}$, we can use the same strategy: let $b_{1}(x)=\frac{1}{|x|}\cdot\mathbf{1}_{\{|x|<1\}}$, where $\mathbf{1}_{A}(x)=\left\{\begin{array}[]{ll}1,&x\in A\\ 0,&x\notin A\end{array}\right$. is the indicator function of $A$. $b_{2}(x)=\frac{1}{|x|}\cdot\mathbf{1}_{\{|x|\geq 1\}}$. It is easy to see $b_{1}\in L^{p}\left(\mathbb{R}^{3}\right)$ for $1\leq p<3$, $b_{2}\in L^{p}\left(\mathbb{R}^{3}\right)$ for $3<p\leq\infty$, and $\frac{1}{|x|}=b_{1}(x)+b_{2}(x)$. Therefore,
$\left\|V_{\rho}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ $=\left\|\frac{1}{|\cdot|}*\rho\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$
$\leq\left\|b_{1}*\rho\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}+\left\|b_{2}*\rho\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$
$\leq\|\rho\|_{L^{p-\epsilon}}\cdot\left\|b_{1}\right\|_{L^{(p-\epsilon)^{\prime}}}+\|\rho\|_{L^{1+\epsilon}}\cdot\left\|b_{2}\right\|_{L^{(1+\epsilon)^{\prime}}}$
The last inequality comes from Young’s Inequality [9, Theorem 4.33 and Exercise 4.30] with $\frac{1}{p-\epsilon}+\frac{1}{(p-\epsilon)^{\prime}}=1$ and $\frac{1}{1+\epsilon}+$ $\frac{1}{(1+\epsilon)^{\prime}}=1$. Since $p>\frac{3}{2}$, we can choose a small $\epsilon$ such that those norms are finite. And then apply the Interpolation Inequality Proposition A.6 for $\|\rho\|_{L^{p-\epsilon}}$ amd $\|\rho\|_{L^{1+\epsilon}}$ to obtain (34). Moreover, we can check $\forall x\in\mathbb{R}^{3},V_{\rho}(x)$ is well defined and finite. (The idea is similar as the following proof of continuity.) Let $\tau_{h}f(x)=f(x+h)$, then
$\left|V_{\rho}(x+h)-V_{\rho}(x)\right|=$ $\left|\int\frac{\rho(x+h-y)-\rho(x-y)}{|y|}d^{3}y\right|$
$\leq$ $\left|\int\left(\tau_{h}\rho(x-y)-\rho(x-y)\right)b_{1}(y)d^{3}y\right|$
$+\left|\int\left(\tau_{h}\rho(x-y)-\rho(x-y)\right)b_{2}(y)d^{3}y\right|$
$\leq$ $\left\|\tau_{h}\rho-\rho\right\|_{L^{p}}\cdot\left\|b_{1}\right\|_{L^{p^{\prime}}}+\left\|\tau_{h}\rho-\rho\right\|_{L^{1}}\cdot\left\|b_{2}\right\|_{L^{\infty}}$
The last inequality comes from the Hölder’s inequality. Notice $p>\frac{3}{2}$ thus its Hölder conjugate $p^{\prime}<3$ thus $\left\|b_{1}\right\|_{L^{p^{\prime}}}<\infty$. $\left\|\tau_{h}\rho-\rho\right\|_{L^{p}}$ and $\left\|\tau_{h}\rho-\rho\right\|_{L^{1}}$ go to 0 when $h\rightarrow 0$. Therefore, $V_{\rho}$ is continuous and $\left\|V_{\rho}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ is finite. Thus $V_{\rho}$ is bounded, i.e. $\exists C>0,\forall x\in\mathbb{R}^{3},|V_{\rho}(X)|<C$. ∎
Notice in the proof above we replace “almost everywhere” $\left(\left\|V_{\rho}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}<C\right)$ by “everywhere” $\left(|V_{\rho}(x)|<C\right)$. The following lemma gives another statement which connects weak derivative and classical derivative, which is a supplement to Morrey Inequality (see [9, Theorem 9.12] or [15, Subsection 5.6.2]).
Lemma A.10 (Continuous Differentiable Representative of Potential).
If $V\in W^{1,\infty}\left(\mathbb{R}^{3}\right)$, and its weak derivative and itself are continuous. Then $V$ has a representative which is continuously differentiable and weak derivative coincides with the classical one a.e. in $\mathbb{R}^{3}$.
Proof.
$V\in W^{1,\infty}\left(\mathbb{R}^{3}\right)$. Therefore, $\forall R>0,V$ has a representative which is Lipschitz in $B_{R}(0)$ [15, Subsection 5.8.2], and then for almost every $x\in\mathbb{R}^{3},V$ is differentiable and the weak derivative coincides with the classical one by Rademacher’s theorem [15, Subsection 5.8.3]. To replace “almost everywhere” by “everywhere”, we modify the discussion about Sobolev space in one dimension in [9, Section 8.2]. Without loss of generality, we assume $j=1$, and consider $\widetilde{f}\left(x_{1},x_{2},x_{3}\right)=\int_{0}^{x_{1}}\frac{\partial V}{\partial x_{1}}\left(y,x_{2},x_{3}\right)dy$, where $\frac{\partial V}{\partial x_{1}}$ is continuous by assumption, one can show its weak and classical derivative with respect to $x_{1}$ are $\frac{\partial V}{\partial x_{1}}$ both in $\mathbb{R}^{3}$ and in $\mathbb{R}$. Therefore, $g:=\widetilde{f}-V$ is continuous and has 0 as its weak partial derivative. Consider $g_{n}=\rho_{n}*g$, where $\left\{\rho_{n}\right\}$ is a sequence of mollifiers, the classical derivative $\frac{\partial g_{n}}{\partial x_{1}}=\frac{\partial\rho_{n}}{\partial x_{1}}*g$ [9, Proposition 4.20], while by the definition of weak derivative we have $\frac{\partial\rho_{n}}{\partial x_{1}}*g=\int\frac{\partial\rho_{n}}{\partial x_{1}}(x-y)g(y)dy=\int\rho_{n}(x-y)\frac{\partial g}{\partial x_{1}}(y)dy=0$. Therefore, for all point in $\mathbb{R}^{3},g_{n}\left(x_{1},x_{2},x_{3}\right)=c_{n}\left(x_{2},x_{3}\right)$. Because $g_{n}\rightarrow g$ uniformly on compact sets of $\mathbb{R}^{3}$ [9, Proposition 4.21], in particular converges everywhere, $g\left(x_{1},x_{2},x_{3}\right)=$ $c\left(x_{2},x_{3}\right)$ for some function $c$. Therefore, $\widetilde{f}\left(x_{1},x_{2},x_{3}\right)=c\left(x_{2},x_{3}\right)+V_{\rho}\left(x_{1},x_{2},x_{3}\right)$ everywhere in $\mathbb{R}^{3}$. Then we take the classical derivative with respect to $x_{1}$ again to see $V_{\rho}$’s classical derivative is $\frac{\partial V_{\rho}}{\partial x_{1}}$, which is also continuous. ∎
It can be shown in the distribution sense we have $\frac{\partial V_{\rho}}{\partial x_{j}}(x)=-\int\frac{y_{j}\rho(x-y)}{|y|^{3}}d^{3}y$. Therefore, combine the arguments and statements as above we have:
Proposition A.11 (Differentiability of Potential [5, Proposition 7]).
If $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{p}\left(\mathbb{R}^{3}\right)$ for some $p>3$, then $V_{\rho}\in$ $W^{1,\infty}\left(\mathbb{R}^{3}\right)$ is continuously differentiable and the weak derivative coincides with the classical one for all $x\in\mathbb{R}^{3}$.
Proof.
Similar to the arguments in proof of Proposition A.9, we know $\frac{\partial V_{\rho}}{\partial x_{j}}\in L^{\infty}\left(\mathbb{R}^{3}\right)\cap C\left(\mathbb{R}^{3}\right)$, and together with Proposition A.9 we have $V_{\rho}\in$ $W^{1,\infty}\left(\mathbb{R}^{3}\right)\cap C\left(\mathbb{R}^{3}\right)$. Therefore, by Lemma A.10 we get the results. ∎
Remark A.12.
One can also prove Proposition A.11 by direct computation.
Acknowledgments
The author is partially supported by the National Science Foundation grant DMS-2308208. This work was primarily carried out during the author’s Master’s studies at the University of Bonn. The author thanks Juan Velázquez (master’s thesis advisor) and Dimitri Cobb (second advisor). Thanks also to Lorenzo Pompili, Shao Liu, Xiaopeng Cheng, Bernhard Kepka, Daniel Sánchez Simón del Pino for discussions, and to Théophile Dolmaire and other instructors. The author is grateful to his parents.